### v1

# Separable homology of graphs and the Whitehead complex

Abstract.
We introduce the Whitehead complex, a one-complex associated to a finite regular cover of the rose and show that it is connected if and only if the fundamental group of the associated cover is generated by its intersection with the set of elements in proper free factors of $\mathbf{F}_{n}$. The Whitehead complex admits an action of $\mathrm{Out}(\mathbf{F}_{n})$ by isometries if the associated cover corresponds to a characteristic subgroup of $\mathbf{F}_{n}$. We prove that the Whitehead complex of the rose has infinite diameter and is nonhyperbolic, implying it is not quasi-isometric to the free splitting complex or the free factor complex.
1. Introduction
Let $\Sigma=\Sigma_{g,n}$ be an oriented surface of finite type, $G$ an abelian group, and $\mathcal{C}$ a collection of closed curves on $\Sigma$. For a curve $\gamma\in\Sigma$, we denote by $[\gamma]$ the associated element of $\mathrm{H}_{1}(\Sigma;G)$. For a cover $\pi:\widetilde{\Sigma}\rightarrow\Sigma$, let $\mathrm{H}_{1}^{\mathcal{C}}(\widetilde{\Sigma};G)$ be the subspace of $\mathrm{H}_{1}(\widetilde{\Sigma};G)$ spanned by the following set:
$\{[\gamma]\in\mathrm{H}_{1}(\widetilde{\Sigma};G)\mid\gamma\text{ is a closed curve in }\widetilde{\Sigma},\pi(\gamma)\in\mathcal{C}\}.$
There has been significant recent interest (see [KS16], [FH16], [MP19], [LRSST20], [BPS23], [Klu23]) in questions of the following form: for a collection of curves $\mathcal{C}\subset\pi_{1}(\Sigma)$, does $\mathrm{H}_{1}^{\mathcal{C}}(\widetilde{\Sigma};G)=\mathrm{H}_{1}(\widetilde{\Sigma};G)$ for all finite regular covers $\widetilde{\Sigma}$ of $\Sigma?$ Since the fundamental group of a punctured surface is free, there are analogous questions regarding the homology of finite regular covers of graphs. The primary focus of this paper is the following closely related question.
Question 1.1.
Let $\mathbf{F}_{n}$ be the free group on $n$ generators, $\mathbf{R}_{n}$ the wedge of $n$ circles, and $\mathcal{C}^{\mathrm{sep}}$ the set of words in $\mathbf{F}_{n}$ contained in proper free factors of $\mathbf{F}_{n}$. Is $\pi_{1}(\Gamma)$ generated by $\mathcal{C}^{\mathrm{sep}}\cap\pi_{1}(\Gamma)$ for every finite regular cover $\Gamma\rightarrow\mathbf{R}_{n}?$
Since finite regular covers of the rose correspond to finite-index normal subgroups of $\mathbf{F}_{n}$, the question above is equivalent to the following question: given a finite-index normal subgroup $N<\mathbf{F}_{n}$, is $N$ always generated by $N\cap\mathcal{C}^{\mathrm{sep}}?$
A positive answer to 1.1 would imply that the integral homology of a finite regular cover of $\mathbf{R}_{n}$ is generated by $\mathcal{C}^{\mathrm{sep}},$ as $\mathrm{H}_{1}(\Gamma;\mathbb{Z})$ is the abelianization of $\pi_{1}(\Gamma)$. It would also imply that $\mathrm{H}_{1}^{\mathrm{sep}}(\widetilde{\Sigma};\mathbb{Z})=\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Z})$ for $\widetilde{\Sigma}$ a finite regular cover of a punctured surface $\Sigma$. Importantly, 1.1 is related to the congruence subgroup problem for mapping class groups. For a statement and history of the congruence subgroup problem, see Chapter 4 of [Far06]. Boggi [Bog06, Ken16] has reduced the congruence subgroup problem for the mapping class group of $\Sigma_{g,n}$ to simple–connectivity of a certain procongruence curve complex associated to $\Sigma_{g,n}$. Using her work in [KLS09], Kent can show that, if $\pi_{1}(\widetilde{\Sigma}_{0,5})$ is generated by the nonfilling elements of $\pi_{1}(\Sigma_{0,5})$, then the procongruence curve complex of $\Sigma_{2,0}$ is simply–connected, giving a conditional solution to the congruence subgroup problem in genus two [Ken]. [^1] This has been announced by Boggi [Bog], but the complete proof has not appeared. The set $\mathcal{C}^{\mathrm{sep}}$ properly contains the set of nonfilling curves on a punctured surface, so a positive answer to 1.1 does not imply the congruence subgroup problem for the mapping class group of the closed surface of genus two. However, it would be a partial result toward this goal.
We provide a reformulation of 1.1 in terms of the connectivity of spaces $\mathrm{Wh}(\Gamma)$ naturally associated to finite regular covers $\Gamma\rightarrow\mathbf{R}_{n}$. Let $\Gamma^{\mathrm{sep}}$ be the cover of $\mathbf{R}_{n}$ whose fundamental group is generated by $\mathcal{C}^{\mathrm{sep}}\cap\pi_{1}(\Gamma)$.
Theorem 1.1.
Given a finite-index regular cover $\Gamma$ of $\mathbf{R}_{n}$, $\mathrm{Wh}(\Gamma)$ is connected if and only if $\pi_{1}(\Gamma)=\pi_{1}(\Gamma^{\mathrm{sep}})$.
Thus, if $\mathrm{Wh}(\Gamma)$ is connected for all finite regular covers $\Gamma,$ the answer to 1.1 is yes. Note that we are identifying the fundamental groups of $\Gamma$ and $\Gamma^{\mathrm{sep}}$ with their images inside $\mathbf{F}_{n}$. If $\Gamma$ represents a characteristic subgroup of $\mathbf{F}_{n}$, the complex $\mathrm{Wh}(\Gamma)$ admits a natural action of $\mathrm{Out}(\mathbf{F}_{n})$ by isometries (see 3.6). Each $\mathrm{Wh}(\Gamma)$ is contained in $\mathrm{Wh}(\mathbf{R}_{n})$, which is connected and admits an isometric action of $\mathrm{Out}(\mathbf{F}_{n})$ (see 3.4). Much of this paper discusses the topological and coarse geometric properties of $\mathrm{Wh}(\Gamma)$.
Theorem 1.2.
Let $\Gamma$ be a finite regular cover of $\mathbf{R}_{n}$. The diameter of every component of $\mathrm{Wh}(\Gamma)$ is infinite.
Let $\mathcal{C}^{\mathrm{prim}}$ be the set of all primitive elements of $\mathbf{F}_{n}$. Since $\mathrm{Wh}(\mathbf{R}_{n})$ is a quotient of $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$, and $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ is quasi-isometric to $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$, we obtain the following corollary.
Corollary 1.3.
The graphs $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ and $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$ have infinite diameter.
We also show that $\mathrm{Wh}(\mathbf{R}_{n})$ is not quasi-isometric to many of the other spaces that admit an isometric action of $\mathrm{Out}(\mathbf{F}_{n})$, including the free factor complex and the free splitting complex, because $\mathrm{Wh}(\mathbf{R}_{n})$ is nonhyperbolic.
Theorem 1.4.
$\mathrm{Wh}(\mathbf{R}_{n})$ is nonhyperbolic.
In addition, we obtain the following corollary.
Corollary 1.5.
The Cayley graphs $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ and $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$ are nonhyperbolic.
In [FH16], Farb and Hensel consider a finite graph $X$ and the set of primitive elements of $\mathbf{F}_{n}$. They ask if $\mathrm{H}_{1}^{\mathrm{prim}}(\Gamma;\mathbb{C})=\mathrm{H}_{1}(\Gamma;\mathbb{C})$ for every finite regular cover $\pi:\Gamma\rightarrow X$. They prove that, if the deck group of $\pi$ is abelian or two-step nilpotent, the answer is yes. However, in general, the answer is no. Malestein and Putman state in Theorem C of [MP19] that if $n\geq 2$ and $\mathcal{C}$ is contained in the union of finitely many $\mathrm{Aut}(\mathbf{F}_{n})$-orbits, there are finite-index normal subgroups $R\triangleleft\mathbf{F}_{n}$ such that $\mathrm{H}_{1}^{\mathcal{C}}(R;\mathbb{Q})\neq\mathrm{H}_{1}(R;\mathbb{Q}).$ Since the set of all primitive elements of $\mathbf{F}_{n}$ is itself an $\mathrm{Aut}(\mathbf{F}_{n})$-orbit, this implies that the primitive homology of a regular cover $\Gamma$ of $\mathbf{R}_{n}$ is not necessarily finite-index in the full homology of $\Gamma$. Later, using GAP  [^2] Groups, Algorithms, Programming - a System for Computational Discrete Algebra, Lee, Rosenblum-Sellers, Safin, and Tenie [LRSST20] found covers $\Gamma$ of relatively low degree compared to those of [MP19] which satisfy $\mathrm{H}_{1}^{\mathrm{prim}}(\Gamma;\mathbb{C})\neq\mathrm{H}_{1}(\Gamma;\mathbb{C}).$ The deck groups of these covers are rank-$3$ $p$-groups for $p=2$ of order at most 512.
Let $\mathcal{C}^{\mathrm{scc}}$ be the set of simple closed curves on a surface $\Sigma$ with genus $g$ and $n$ punctures (possibly zero). For $\Sigma$ with large enough complexity, i.e. $(g,n)\not\in\{(0,0),(0,1),(0,2),(1,0)\}$, Koberda and Santharoubane [KS16] give examples of finite covers $\widetilde{\Sigma}\rightarrow\Sigma$ satisfying
$\mathrm{H}_{1}^{\mathrm{scc}}(\widetilde{\Sigma};\mathbb{Z})\neq\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Z})$
using TQFT representations. Furthermore, Theorem C of [MP19] implies that, if $\Sigma$ is a punctured surface of genus at least two, there are finite regular covers $\widetilde{\Sigma}$ of $\Sigma$ satisfying $\mathrm{H}_{1}^{\mathrm{scc}}(\widetilde{\Sigma};\mathbb{Q})\neq\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Q})$. This is because $\mathcal{C}^{\mathrm{scc}}$ is contained in finitely many $\mathrm{Aut}(\mathbf{F}_{n})$-orbits. The same is true for the set of curves on $\Sigma$ with at most $k$ self-intersections. Recently, Klukowski ([Klu23]) showed that for any surface $\Sigma=\Sigma_{g,n}$ satisfying $g\geq 2$ and $n\geq 0$, there is a finite cover $\widetilde{\Sigma}$ of $\Sigma$ such that $\mathrm{H}_{1}^{\mathrm{scc}}(\widetilde{\Sigma};\mathbb{Q})\neq\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Q}).$
In 2008, while working on the congruence subgroup problem for $\mathrm{Mod}(\Sigma_{g})$, Kent asked the following question, which remains open despite the closely related work described above.
Question 1.2.
[Ken12] Let $\Sigma$ be a compact oriented surface, $\mathcal{C}^{\mathrm{nf}}$ the set of all nonfilling curves on $\Sigma$, and $\pi:\widetilde{\Sigma}\rightarrow\Sigma$ a finite regular cover. Is $\mathrm{H}_{1}^{\mathrm{nf}}(\widetilde{\Sigma};\mathbb{Z})=\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Z})?$
Using Boggi’s original approach to the congruence subgroup problem (see [Bog06]) and the ideas of Kent mentioned previously, a positive answer to this question would also give a proof of the congruence subgroup problem in genus two. Recently, Boggi, Putman, and Salter [BPS23] made significant progress on 1.2: they showed that there is a set of nonfilling curves on $\Sigma$–namely, the set $\mathcal{C}^{\mathrm{pant}}$ of curves contained in pairs of pants–which generates the rational homology of every finite branched cover of a closed oriented surface. This does not fully answer 1.2, but it does imply that $\mathrm{H}_{1}^{\mathrm{nf}}(\widetilde{\Sigma};\mathbb{Z})$ is finite-index in $\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Z})$ for any finite cover $\widetilde{\Sigma}$ of a closed oriented surface $\Sigma$.
We say an element of $\mathbf{F}_{n}$ is separable if it is contained in a proper free factor of $\mathbf{F}_{n}$. Note that the set of nonfilling curves on a punctured surface $\Sigma$ is contained in the set of separable curves on $\Sigma$. Therefore, if $\mathrm{H}_{1}^{\mathrm{sep}}(\Gamma;\mathbb{Z})\neq\mathrm{H}_{1}(\Gamma;\mathbb{Z}),$ $\mathrm{H}_{1}^{\mathrm{nf}}(\widetilde{\Sigma};\mathbb{Z})\neq\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Z}).$ A weaker version of 1.1 is below.
Question 1.3.
Is $\mathrm{H}_{1}(\Gamma;\mathbb{Z})$ generated by $\mathcal{C}^{\mathrm{sep}}$ for every finite regular cover $\Gamma\rightarrow\mathbf{R}_{n}?$
We give the following reformulation of this question.
Theorem 1.6.
Given a finite-index regular cover $\Gamma$ of $\mathbf{R}_{n}$, there is a naturally defined space $\mathrm{W}_{1}(\Gamma)$ which is connected if and only if $\mathrm{H}_{1}^{\mathrm{sep}}(\Gamma;\mathbb{Z})=\mathrm{H}_{1}(\Gamma;\mathbb{Z}).$
Finally, using the result of Boggi, Putman, and Salter mentioned above, we obtain the following.
Theorem 1.7.
Let $\Gamma\rightarrow\mathbf{R}_{n}$ be a finite regular cover. Then $\mathrm{W}_{1}(\Gamma)$ has finitely many components.
Outline
In Section 2, we describe the connection between graphs and free groups (Section 2.1), state some important facts about Van Kampen diagrams (Section 2.2), and recall an algorithm of Whitehead which can determine whether $w\in\mathbf{F}_{n}$ is separable (Section 2.3). In Section 3, we define the Whitehead complex (Definition 3.1) and prove 1.1. In Section 3.1 we prove 1.2 and 1.3. We dedicate Section 4 to the proof of 1.4 and 1.5. Finally, Section 5 contains a proof of 1.6 and 1.7.
Acknowledgments
The author would like to thank her advisors, Autumn Kent and Marissa Loving, for their invaluable help on this project. She would also like to thank Caglar Uyanik for many helpful conversations. She is grateful to Alex Wright for his comments on 1.5. In particular, Wright pointed out that $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ and $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$ are quasi-isometric. She also thanks Adam Klukowski, Jean Pierre Mutanguha, Nick Salter, and Alex Wright for helpful comments on a preprint of this paper. This work is supported in part by NSF Grants DMS-2202718, RTG DMS-2230900, and DMS-1904130.
2. Background
Here we cover some prerequisite information for the rest of the paper. There is information about graphs and free groups in Section 2.1. In Section 2.2, we discuss Van Kampen diagrams and annular diagrams, which are essential tools in Section 3 and Section 4. In Section 2.3, we discuss an algorithm of Whitehead which decides whether $w\in\mathbf{F}_{n}$ is in a proper free factor of $\mathbf{F}_{n}.$
2.1. Graphs and Free Groups
Definition 2.1.
A directed graph is a 1-complex given by a tuple $(V,E,\iota,\tau)$ where $V$ is a set of 0-cells called vertices, $E$ is a set of 1-cells called edges, and $\iota:E\rightarrow V$ and $\tau:E\rightarrow V$ are maps that send $e\in E$ to the endpoints of $e$ in $V$. For all $e\in E$, $\iota(e)$ will be called the initial vertex of $e$ and $\tau(e)$ will be called the terminal vertex of $e$. By assigning initial and terminal vertices of $e$, we assign an orientation to $e$. We represent the orientation of $e$ by an arrow on $e$ pointing toward $\tau(e)$. The valence of a vertex $v$ is the number of components of $N(v)\setminus\{v\}$, where $N(v)$ is a small neighborhood of $v$.
Note that, under our definition of graph, we allow multiple edges with the same initial and terminal vertices. Often we will want to label the edges or vertices with elements of some set. We will denote the label of an edge $e$ or vertex $v$ by $\mathrm{lab}(e)$ and $\mathrm{lab}(v)$, respectively.
Definition 2.2.
A rose is a graph with one vertex and one or more edges. We denote the rose with $n$ edges by $\mathbf{R}_{n}$.
Throughout the paper, we let $\mathbf{F}_{n}$ be the free group on $n$ generators
$X_{n}:=\{x_{1},x_{2},\ldots,x_{n}\}.$
We denote the inverse of an element $v\in\mathbf{F}_{n}$ by $\overline{v}$. Let $X_{n}^{*}:=\{x_{1},x_{2},\ldots,x_{n},\overline{x_{1}},\overline{x_{2}},\ldots,\overline{x_{n}}\}$. A word on $X_{n}^{*}$ is an element of the set
$\mathcal{W}_{n}=\{w\in\displaystyle\prod_{i=1}^{m}X_{n}^{*}\mid m<\infty\}.$
Elements of $\mathbf{F}_{n}$ are equivalence classes of words under free reduction. A word is freely reduced if it contains no subword of the form $y\overline{y}$ where $y\in X_{n}^{*}$. The freely reduced word $w=y_{1}y_{2}\ldots y_{m}$ where $y_{i}\in X^{*}$ is cyclically reduced if $y_{1}\neq\overline{y_{m}}$. A cyclic subword of a word $w$ is a subword of one of the cyclic permutations of $w$. The length of $w=y_{1}y_{2}\ldots y_{m}$ where $y_{i}\in X_{n}^{*}$ is denoted by $|w|$: in this case, $|w|=m$. When we refer to the length of $v\in\mathbf{F}_{n}$, we mean the length of the freely reduced word representing $v$. For an automorphism $\varphi:\mathbf{F}_{n}\rightarrow\mathbf{F}_{n}$, there is a unique map $\tilde{\varphi}:\mathcal{W}_{n}\rightarrow\mathcal{W}_{n}$ which extends $\varphi$: for every $y\in X_{n}^{*}$, let $\tilde{\varphi}(y)=\varphi(y).$
Since $\pi_{1}(\mathbf{R}_{n})\cong\mathbf{F}_{n}$, we will assume throughout that the edges of $\mathbf{R}_{n}$ are labeled by elements of $X_{n}$. A cover of $\mathbf{R}_{n}$ will inherit this labeling, so that a graph $\Gamma$ is a cover of $\mathbf{R}_{n}$ if and only if for every vertex $v$ in $\Gamma$, there is a neighborhood $U$ of $v$ consisting of $v$ and two edge segments incident to $v$ for each $x_{i}\in X_{n}$, oriented in opposite directions. We identify $\pi_{1}(\mathbf{R}_{n})$ with $\mathbf{F}_{n}$ and, for $v$ a vertex of $\Gamma$, we identify $\pi_{1}(\Gamma,v)$ with the corresponding subgroup of $\mathbf{F}_{n}$.
Every finite-index subgroup of $\mathbf{F}_{n}$ is represented by a finite cover of $\mathbf{R}_{n}$. Suppose $G<\mathbf{F}_{n}$ is a finitely generated subgroup of $\mathbf{F}_{n}$ that is not finite-index. Then there is a finite graph called a core graph such that every element of $G$ is represented by a closed path in $G$ with no backtracking. In order to define a core graph we need to introduce the notion of folding.
Definition 2.3.
A Stallings folding, or simply folding, of a labeled, directed graph is an identification of edges $e_{1},e_{2}$ that satisfy the following:
  1. (1)
    $\iota(e_{1})=\iota(e_{2})$ or $\tau(e_{1})=\tau(e_{2});$
  2. (2)
    $\mathrm{lab}(e_{1})=\mathrm{lab}(e_{2});$
  3. (3)
    $e_{1}$ and $e_{2}$ have the same orientation.
If a labeled, directed graph has no pairs of edges satisfying the conditions above, we say the graph is completely folded.
See Figure 1 for an example of a folding. Note that foldings induce surjections on the corresponding fundamental groups: see [Sta83, Corollary 4.4].
Figure 1. A Folding
Definition 2.4.
A core graph is a finite completely folded graph with at most one valence-one vertex.
To obtain a core graph representing a finitely-generated subgroup $G=\langle w_{1},w_{2},\ldots,w_{k}\rangle$ of $\mathbf{F}_{n}$, take the wedge sum of the $k$ subdivided cycles labeled by the generators of $G$, then completely fold. The wedge point is the basepoint of the graph. See [Sta83, Section 5.4] for a proof of this fact. An example of obtaining a core graph via complete folding is shown in Figure 2 for $G=\langle ab\overline{a},[a,b]\rangle$, where $[a,b]=ab\overline{a}\overline{b}$. We will use this notation for commutators throughout the paper.
Figure 2. Folding to obtain a core graph for $G=\langle a,b\mid ab\overline{a},[a,b]\rangle$
We will introduce the Whitehead complex of the rose $\mathrm{Wh}(\mathbf{R}_{n})$ in Section 3 and prove that it admits an isometric action of $\mathrm{Out}(\mathbf{F}_{n})$. Thus, we include here a brief discussion of $\mathrm{Out}(\mathbf{F}_{n})$ and some of the spaces on which it acts. $\mathrm{Aut}(\mathbf{F}_{n})$ is the group of automorphisms of the free group. A theorem of Nielsen (see [MKS04, Theorem 3.2]) gives a nice set of generators for $\mathrm{Aut}(\mathbf{F}_{n}).$
Theorem 2.5 (Nielsen).
Let $X_{n}=\{x_{1},\ldots,x_{n}\}$ be a basis for $\mathbf{F}_{n}$. $\mathrm{Aut}(\mathbf{F}_{n})$ is finitely generated by the following elementary Nielsen transformations.
  1. (1)
    Let $\sigma$ be a permutation of $\{1,2,\ldots,n\}$ and $\epsilon_{i}\in\{\pm 1\}$ for $1\leq i\leq n$. A function $N_{\sigma}:\mathbf{F}_{n}\rightarrow\mathbf{F}_{n}$ satisfying
    $N_{\sigma}(x_{i})=x_{\sigma(i)}^{\epsilon_{i}}$
    for all $i$ is an elementary Nielsen transformation.
  2. (2)
    Let $i$ and $j$ satisfy $1\leq i,j\leq n$. A function $N:\mathbf{F}_{n}\rightarrow\mathbf{F}_{n}$ fixing all $x_{k}$ for $k\neq i$ and sending $x_{i}$ to an element of $\{x_{i}x_{j}^{\pm 1},x_{j}^{\pm 1}x_{i},x_{j}^{\epsilon}x_{i}x_{j}^{-\epsilon}\}$ where $\epsilon\in\{\pm 1\}$ is also an elementary Nielsen transformation.
Remark 2.6.
A straightforward corollary of this theorem is that if the multiset of letters of $w\in\mathbf{F}_{n}$ contains only one element of $\{x,\overline{x}\}$ for some $x\in X_{n}$, then $w$ is primitive. To see this, observe that there is a composition of Nielsen automorphisms taking $x$ to $xu,$ where $u\in\mathbf{F}\langle X_{n}\setminus\{x\}\rangle.$
For $n\geq 2$, $\mathrm{Out}(\mathbf{F}_{n})$ acts on the free factor complex $\mathrm{FF}_{n}$ (first introduced by Hatcher and Vogtmann in [HV98]) and its 1-skeleton, the free factor graph. A theorem of Bestvina and Feign [BF14] states that for $n\geq 2$, $\mathrm{FF}_{n}$ is Gromov-hyperbolic (see also [KR14, Theorem 6.1] and [HH17]). In addition, $\mathrm{Out}(\mathbf{F}_{n})$ acts on the free bases complex, denoted $\mathrm{FB}_{n}$; by [KR14, Proposition 4.3], $\mathrm{FB}_{n}$ is quasi-isometric to $\mathrm{FF}_{n}.$ $\mathrm{Out}(\mathbf{F}_{n})$ also acts on the free splitting complex, which has a Lipschitz map to the free factor graph [KR14, Remark 4.5]. The free splitting complex was first shown to be hyperbolic by Handel and Mosher in [HM13].
2.2. Van Kampen diagrams
Throughout the paper, we make use of Van Kampen diagrams, first introduced in 1933 by E. R. Van Kampen. For a more complete discussion, see Chapter V of [LSLS77]. Let $\mathcal{P}=\langle X\mid R\rangle$ be group presentation. We will assume throughout the paper that elements of $R$ are cyclically reduced. Let $R^{*}$ be the symmetric closure of $R$: that is, the set of cyclic permutations of elements of $R$ and their inverses.
Definition 2.7.
Let $\mathcal{P}=\langle X\mid R\rangle$ be a presentation for a group, and assume elements of $R$ are cyclically reduced. A Van Kampen diagram for $w\in\langle\langle R\rangle\rangle$ over $\mathcal{P}$ is a finite, simply connected planar 2-complex $D$ with oriented edges labeled by elements of $X$. The boundary of $D$ is labeled by $w$ beginning at some basepoint for $D$. The 1-skeleton of $D$ is a folding (not necessarily a complete folding) of a wedge of cycles labeled by conjugates of elements of $R^{*}$. The 2-disks of $D$ are called regions. Regions are attached along cycles labeled by elements of $R^{*}$. Regions of $D$ have completely folded boundaries which are embedded in $D$. The area of $D$ is the number of regions of $D$. A Van Kampen diagram $D$ for $w$ over $\mathcal{P}$ is minimal if there is no Van Kampen diagram $D^{\prime}$ for $w$ over $\mathcal{P}$ of smaller area.
Let $G$ be a quotient of $\mathbf{F}_{n}$, and let $\mathcal{P}=\langle X_{n}\mid R\rangle$ be a presentation for $G$. A word $w\in\langle\langle R\rangle\rangle$ has a Van Kampen diagram over $\mathcal{P}$. The converse is also true: if $w\not\in\langle\langle R\rangle\rangle$, then $w$ has no Van Kampen diagram over $\mathcal{P}$.
Remark 2.8.
When constructing Van Kampen diagrams over a presentation $\langle X\mid R\rangle$, we adopt the convention that the boundaries of disks are glued along cycles labeled by elements of $R^{*}$ (as opposed to conjugates of elements of $R^{*}$). For example, if $w=aba\overline{c}\overline{b^{2}}ab\overline{a}\overline{b}\overline{a}$, the boundary of a region corresponding to $w$ in a Van Kampen diagram will be glued along a cycle labeled by $\overline{c}\overline{b^{2}}ab$, not $w$. The complex obtained by gluing disks along cycles labeled by elements of $R^{*}$ is homotopy equivalent to the complex obtained by gluing disks along cycles labeled by conjugates of elements of $R^{*}$. This will simplify certain arguments in Section 3 and Section 4.
Let $D(w)$ be a Van Kampen diagram over a presentation $\mathcal{P}$ for a word $w$, and let $cw\overline{c}$ be any conjugate of $w$. There is a diagram for $cw\overline{c}$ which differs from $D(w)$ only by a change of basepoint and, possibly, the addition or deletion of a dead-end path beginning at some point on the boundary of $D(w)$.
Definition 2.9.
A diagram over $\langle X\mid R\rangle$ for a conjugacy class $[w]$ is the 2-complex resulting from taking a Van Kampen diagram for any $\hat{w}\in[w]$, forgetting the basepoint, and deleting all valence-one vertices.
We will also utilize annular diagrams in Section 4. Annular diagrams over a presentation $\mathcal{P}=\langle X\mid R\rangle$ are similar to Van Kampen diagrams in that they are planar 2-complexes and their regions have boundaries labeled by elements of $R^{*}$, but they are not simply connected.
Definition 2.10.
Let $\langle X\mid R\rangle$ be a group presentation, and assume that elements of $R$ are cyclically reduced. An annular diagram $A$ over $\langle X\mid R\rangle$ is a finite connected planar 2-complex with fundamental group $\mathbb{Z}$ and two distinct, though not necessarily disjoint, boundaries (see Figure 3). The boundaries of the regions of annular diagrams are labeled by elements of $R^{*}$.
Figure 3. Two annular diagrams with boundaries $w_{1}$ and $w_{2}$. Note that in the right diagram, the two boundaries overlap yet are distinct.
2.3. Whitehead’s algorithm for separable words
A proper free factor of $\mathbf{F}_{n}$ is a nontrivial proper subgroup $F<\mathbf{F}_{n}$ such that $\mathbf{F}_{n}$ splits as a free product $\mathbf{F}_{n}=F*G$ for some nontrivial, proper $G<\mathbf{F}_{n}$. We will need to be able to distinguish between words that are in a proper free factor of $\mathbf{F}_{n}$ and words that are not.
Definition 2.11 ([Sta99]).
A word in $\mathbf{F}_{n}$ is separable if it belongs to a proper free factor of $\mathbf{F}_{n}$. Otherwise, $w$ is inseparable.
Some authors refer to these words as simple; we use Stallings’ terminology in order to avoid confusion with the definition of simple curves on a surface. The proof of the following theorem of Stallings relies on ideas of Whitehead in [Whi36], hence the name “Whitehead’s algorithm for separable words.”
Theorem 2.12 ([Sta99]).
Given a word $w\in\mathbf{F}_{n},$ there is an efficient algorithm to determine whether or not $w$ is separable.
In fact, Algorithm 2.5 of [Sta99] determines whether finite sets of words $S\subset\mathbf{F}_{n}$ are separable (i.e., whether all elements of $S$ are conjugate into the factors of any splitting $F\star G=\mathbf{F}_{n}$). We will only need the result for words in $\mathbf{F}_{n}$, so our definitions and algorithm only refer to separability of a single word. The algorithm to determine separability of $w$ constructs a graph called the Whitehead graph of $w$, denoted $\Omega(w)$. Recall that, for a cyclically reduced word $w$, a cyclic subword of $w$ is a subword of one of the cyclic permutations of $w$.
Definition 2.13 (Whitehead graph of a word).
Let $w\in\mathbf{F}_{n}$, where $\mathbf{F}_{n}$ is generated by $X_{n}$. The Whitehead graph of $w$, denoted $\Omega(w)$, is a directed graph constructed as follows:
  • The vertex set of $\Omega(w)$ is $X_{n}^{*}$;
  • There is one edge from $x$ to $y$ for each distinct copy of the cyclic subword $x\overline{y}$ in $w$; thus the edge set $E$ of $\Gamma$ is in bijection with the multiset of length two cyclic subwords of $w.$
Equivalently, we can define the Whitehead graph as follows. Let $M$ be the connected sum of $n$ copies of $\mathbb{S}^{1}\times\mathbb{S}^{2}$. The fundamental group of $M$ is $\mathbf{F}_{n}$, and the generators of $\pi_{1}(M)$ are represented by a system of $n$ disjoint 2-spheres $\{S_{1},S_{2},\ldots,S_{n}\}\subset M$. Each sphere has a positive side $S_{i}^{+}$ and a negative side $S_{i}^{-}$. The element $x_{i}\in\mathbf{F}_{n}$ is represented by a loop beginning at the basepoint of $M$ which travels to $S_{i}^{-}$, pierces $S_{i}$, and returns to the basepoint from $S_{i}^{+}$. Here we choose the system of spheres so that no sphere intersects the basepoint of $M$. The advantage of considering curves in $M$ instead of $\mathbf{R}_{n}$ is that every $w\in\mathbf{F}_{n}$ can be represented by an embedding of $\mathbb{S}^{1}$ in $M,$ whereas in $\mathbf{R}_{n}$ this is impossible. Cutting $M$ along the sphere system results in a handlebody $\hat{M}$ with $2n$ spheres on its boundary. Let $w\in\mathbf{F}_{n}$, and represent $w$ as a geodesic curve $c\in M$. The Whitehead graph of $w$ is the graph whose edges are the arcs of $c$ in $\hat{M}$ and whose vertices are the $2n$ boundary spheres of $\hat{M}$. The orientations of the edges in the Whitehead graph of $w$ are given by the orientations of the arcs of $c$ in $\hat{M}$. An example of the Whitehead graph for a specific word in $\mathbf{F}_{n}$ is shown in Figure 4.
Figure 4. The Whitehead graph of $ab^{2}\overline{c}da\overline{b}d^{3}\overline{a}\overline{c^{2}}\in\mathbf{F}_{4}$.
Definition 2.14.
A vertex $v$ in a graph $\Gamma$ is a cut vertex if, after removing $v$ and all edges incident to $v$, the resulting graph is disconnected.
Whitehead’s algorithm to decide separability of $w$ begins by constructing $\Omega(w)$. If $\Omega(w)$ is connected with no cut vertices, $w$ is inseparable, and if $\Omega(w)$ is disconnected, $w$ is separable. In the remaining case, $\Omega(w)$ is connected and has a cut vertex $x$. In this case, we apply a particular automorphism to $w$ and reconstruct $\Omega(w)$. The type of automorphisms used in the algorithm are called Whitehead automorphisms of the second kind.
Definition 2.15.
A Whitehead automorphism of the second kind is an element $\phi\in\mathrm{Aut}(\mathbf{F}_{n})$ satisfying the property that there is some fixed $x\in X_{n}^{*}$ called the multiplier such that for $a\in X_{n}^{*},$ $\phi(a)\in\{a,xa,a\overline{x},xa\overline{x}\}$.
Note that Whitehead automorphisms of the first kind are the elements of $\mathrm{Aut}(\mathbf{F}_{n})$ that permute the generators and their inverses. The set of Whitehead automorphisms of the first and second kinds generates $\mathrm{Aut}(\mathbf{F}_{n})$ as it contains the elementary Nielsen transformations.
Given a choice of $x\in X_{n}^{*}$ and a partition $Y\bigsqcup Z$ of $X_{n}^{*}$ into two nonempty subsets such that $x\in Y$ and $\overline{x}\in Z$, we define a Whitehead automorphism $\phi_{x,Y,Z}$ as follows:
$\varphi_{x,Y,Z}(y)=\begin{cases}y&\quad\text{if }x=y\text{ or }\{y,\overline{y}\}\subset Y;\\ xy&\quad\text{if }y\in Y\text{ and }\overline{y}\in Z;\\ y\overline{x}&\quad\text{if }y\in Z\text{ and }\overline{y}\in Y;\\ xy\overline{x}&\quad\text{if }y,\overline{y}\in Z.\\ \end{cases}$
Theorem 2.16 ([Sta99]).
The following process will determine whether $w\in\mathbf{F}_{n}$ is separable in finite time.
  1. (1)
    Construct $\Omega(w)$.
  2. (2)
    If $\Omega(w)$ is disconnected, $w$ is separable. [Sta99, Proposition 2.2]
  3. (3)
    Else if $\Omega(w)$ is connected with no cut vertices, $w$ is inseparable. [Sta99, Theorem 2.4]
  4. (4)
    Else $\Omega(w)$ is connected and has a cut vertex $x\in X_{n}^{*}$. Partition $X_{n}^{*}$ into two disjoint subsets $Y,Z$ as follows: let $\Omega^{\prime}$ be the graph resulting from the removal of $x$ and all edges incident to $x$. Let $Z$ be the vertices in the component of $\Omega^{\prime}$ containing $\overline{x}$, and let $Y=X_{n}^{*}\setminus Z$. Return to Step 1 and repeat the process for $\phi_{x,Y,Z}(w);$ eventually, this process terminates [Sta99, Proposition 2.3].
A proof that this algorithm works can be found in [Sta99].
Example 2.17.
We will use Whitehead’s algorithm to show $w=a\overline{b}cb\overline{a}\,\overline{c}a$ is separable in $\mathbf{F}\langle a,b,c\rangle.$ See Figure 5. Observe that $a$ is a cut vertex. The partition of the vertices is given by $\{a,b\}\bigsqcup\{\overline{a},\overline{b},c,\overline{c}\}.$ The automorphism associated with this partition is below.
$\phi(x)=\begin{cases}a&x=a\\ ab&x=b\\ ac\overline{a}&x=c\end{cases}$
Apply $\phi$ to $w$ to obtain $\phi(w)=a\overline{b}cb\overline{c}.$ Observe that $w$ is primitive as it contains only one copy of $a$ or its inverse, so $w$ is separable.
Figure 5. On the left is the Whitehead graph for $w=a\overline{b}cb\overline{a}\,\overline{c}a$; on the right, the Whitehead graph for $\phi(w)=a\overline{b}cb\overline{c}$.
Definition 2.18.
The profile of a Whitehead graph $\Omega$ is an undirected finite graph $\sigma(\Omega)$ which is the image of $\Omega$ under the graph surjection $\sigma$ which takes all directed edges of $\Omega$ between vertices $x$ and $y$ to a single undirected edge between $x$ and $y$.
See Figure 6 for an example of a Whitehead profile. The advantage of considering the profile of a Whitehead graph is that there are finitely many profiles of Whitehead graphs of curves $\gamma$ in $\Gamma$, and one can determine from the Whitehead profile whether or not the Whitehead graph has a cut vertex. Showing that the Whitehead profile of $w\in\mathbf{F}_{n}$ has no cut vertex is a quick way to ensure $w$ is inseparable.
Figure 6. On the left, the Whitehead graph of $a^{2}bc\overline{d}a\overline{d}abc\overline{b}\overline{a}d^{2}$; on the right, its Whitehead profile
Another variation of the Whitehead graph that is sometimes useful is the labeled Whitehead graph, defined below.
Definition 2.19.
The labeled Whitehead graph of a cyclically reduced $w\in\mathbf{F}_{n}$ is the Whitehead graph $\Omega(w)$ together with a labeling of the edges of $\Omega(w)$ by $\mathbb{Z}\big{/}m\mathbb{Z}$, where $m$ is the length of $w$. The edges of $\Omega(w)$ are labeled in the order that the corresponding length-two cyclic subwords of $w$ appear in $w$. Two labeled Whitehead graphs $\Omega$ and $\Psi$ are equivalent whenever $\Omega$ and $\Psi$ are isometric via an isometry that preserves vertex labels, and the edge labels of $\Omega$ and $\Psi$ differ only by a fixed element of $\mathbb{Z}\big{/}m\mathbb{Z}$. See Figure 7 for an example.
Equivalence classes of labeled Whitehead graphs are in one-to-one correspondence with conjugacy classes of elements of the free group.
Figure 7. Equivalent labeled Whitehead graphs of $ab^{2}\overline{c}da\overline{b}d^{3}\overline{a}\overline{c^{2}}$
3. The Whitehead complex
The purpose of this section is to introduce the Whitehead complex and describe some of its properties. The Whitehead complex is a tool to study finite regular covers $\Gamma$ of the rose (equivalently, finite-index normal subgroups of $\mathbf{F}_{n}$). A key property of $\mathrm{Wh}(\Gamma)$ is that it is connected if and only if $\pi_{1}(\Gamma)$ is generated by $\mathcal{C}^{\mathrm{sep}}\cap\pi_{1}(\Gamma)$ (see 1.1). Throughout the paper, $\Gamma$ is a finite-index regular cover of $\mathbf{R}_{n}$ for $n\geq 2$, and $\Gamma^{\mathrm{sep}}$ is the cover of $\mathbf{R}_{n}$ representing the subgroup of $\pi_{1}(\Gamma)$ generated by elements of $\mathcal{C}^{\mathrm{sep}}\cap\pi_{1}(\Gamma)$. See Figure 8 for a small subgraph of the Whitehead complex of a specific cover of $\mathbf{R}_{n}$.
Definition 3.1 (Whitehead complex).
Let $\Gamma$ be a finite regular cover of $\mathbf{R}_{n}$ for some $n\geq 2$. The Whitehead complex of $\Gamma$, denoted $\mathrm{Wh}(\Gamma)$, is a locally infinite directed graph constructed as follows:
  • The vertex set $V$ of $\mathrm{Wh}(\Gamma)$ is the set of equivalence classes of elements of $\pi_{1}(\Gamma)$ under conjugation by $\mathbf{F}_{n}$;
  • Each edge is labeled by the (free) conjugacy class of a nontrivial element of $\pi_{1}(\Gamma^{\mathrm{sep}})$. There is an edge $[\alpha]$ from $[v]$ to $[w]$ whenever there are $\hat{\alpha}\in[\alpha],$ $\hat{w}\in[w],$ $\hat{v}\in[v]$ such that $\hat{w}=\hat{v}\hat{\alpha}$ in $\mathbf{F}_{n}$.
We will generally use Latin letters in square brackets to refer to vertices in $\mathrm{Wh}(\Gamma)$ and Greek letters in square brackets to denote edge labels in $\mathrm{Wh}(\Gamma)$. Note that we can take $\hat{w}$ to be any representative of $[w]$ by conjugating the equation $\hat{w}=\hat{v}\hat{\alpha}$.
Remark 3.2.
Equivalently, we can think of vertices of $\mathrm{Wh}(\Gamma)$ as equivalence classes of labeled Whitehead graphs (see Definition 2.19). Two vertices $[\Omega(v)]$ and $[\Omega(w)]$ are adjacent whenever there is an $\alpha\in\mathcal{C}^{\mathrm{sep}}$ and a representative $\hat{v}\in[v]$ such that $\Omega(\hat{v}\alpha)\in[\Omega(w)].$ An edge in $\mathrm{Wh}(\Gamma)$ under this definition of $\mathrm{Wh}(\Gamma)$ corresponds to a labeled Whitehead graph of a separable word. The name “Whitehead complex” comes from this alternate definition of $\mathrm{Wh}(\Gamma)$.
Figure 8. Above, a finite regular cover $\Gamma$ of $\mathbf{R}_{3}$; below, a small portion of $\mathrm{Wh}(\Gamma)$ to illustrate edge relations.
Every vertex in $\mathrm{Wh}(\Gamma)$ is incident to at least one edge for each conjugacy class of nontrivial curves in $\Gamma^{\textrm{sep}}$, so $\mathrm{Wh}(\Gamma)$ is locally infinite. We think of $\mathrm{Wh}(\Gamma)$ as a metric space with the induced path metric: every edge of $\mathrm{Wh}(\Gamma)$ has length one, and the distance between $[v]$ and $[w]\in\mathrm{Wh}(\Gamma)$ is the length of the shortest path between $[v]$ and $[w]$. If no path between $[v]$ and $[w]$ exists, the distance between them is infinite. The distance between $[v]$ and $[w]$ is denoted $d_{\mathrm{sep}}\Big{(}[v],[w]\Big{)}$. The distance between $[1]$ and $[w]\in\mathrm{Wh}(\Gamma)$ is called the separable length of $[w]$ (or the separable length of $w$) and denoted by $||w||_{\mathrm{sep}}$.
Consider the set of paths in $\mathrm{Wh}(\Gamma)$, denoted $\mathcal{P}\mathrm{Wh}(\Gamma)$. Let $p([v],[w])\in\mathcal{P}\mathrm{Wh}(\Gamma)$ have edges labeled by $[\alpha_{1}],[\alpha_{2}],\ldots,[\alpha_{l(p)}]$, where $l(p)$ is the length of $p$. Then $p$ corresponds to an expression for $\hat{w}\in[w]$ as a product of a representative $\hat{v}$ of $[v]$ and elements $\hat{\alpha_{1}},\hat{\alpha_{2}},\ldots\hat{\alpha_{l(p)}}$ of the separable conjugacy classes labeling the edges of $p$. We describe in Lemma 3.3 why the existence of a path from $[1]$ to $[w]$ implies the existence of a Van Kampen diagram for $[w]$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\cap\pi_{1}(\Gamma)\rangle$.
Lemma 3.3.
If $p([1],[w])\in\mathcal{P}\mathrm{Wh}(\Gamma)$, there is a Van Kampen diagram $D_{p}$ for $[w]$ over $\langle X_{n}\mid\pi_{1}(\Gamma^{\mathrm{sep}})\rangle.$ The regions of $D_{p}$ have boundaries labeled by representatives of the conjugacy classes labeling the edges of $p$.
Proof.
Let the edges of $p([1],[w])$ be $[\alpha_{1}],[\alpha_{2}],\ldots,[\alpha_{l}]$ where $l$ is the length of $p([1],[w]).$ By the definition of $\mathrm{Wh}(\Gamma)$, there are $\hat{\alpha_{i}}\in[\alpha_{i}]$ so that $\hat{\alpha_{1}}\hat{\alpha_{2}}\ldots\hat{\alpha_{l}}=\hat{w}\in[w]$. To construct a diagram for $[w]$, wedge disks with the same orientation and boundary labels $\hat{\alpha_{1}},\hat{\alpha_{2}},\ldots,\hat{\alpha_{l}}$ at a point to obtain a 2-complex $D$, then fold until the boundary of the 2-complex is labeled by $\hat{w}$. Delete all valence-one vertices to get a diagram for $[w].$ See Figure 9 for an example of the diagram corresponding to a specific path in the Whitehead complex. ∎
Figure 9. A path in $\mathrm{Wh}(\mathbf{R}_{3})$ from $1$ to $[abc^{2}\overline{a}\overline{b}]$ and a corresponding sequence of Van Kampen diagrams.
See 1.1
Proof.
Each path $p=p([1],[v])\in\mathcal{P}\mathrm{Wh}(\Gamma)$ corresponds to an expression $\prod_{i=1}^{l}\alpha_{i}=\hat{v}\in[v]$, where $l$ is the length of $p$ and $\alpha_{i}\in\pi_{1}(\Gamma^{\mathrm{sep}})$ are representatives of the conjugacy classes labeling the edges of $p$. If $\mathrm{Wh}(\Gamma)$ is connected, there is a path from $[1]$ to every vertex $[v]\in\mathrm{Wh}(\Gamma)$. Since $\Gamma$ is regular, the existence of an expression for $\hat{v}$ as a product of separable elements implies that any $v^{\prime}\in[v]$ is expressible as a product of elements of $\pi_{1}(\Gamma^{\mathrm{sep}})$. Hence $\pi_{1}(\Gamma)=\pi_{1}(\Gamma^{\mathrm{sep}})$. Similarly, supposing $\pi_{1}(\Gamma)=\pi_{1}(\Gamma^{\mathrm{sep}}),$ there are expressions for all $g\in\pi_{1}(\Gamma)$ as products of elements of $\pi_{1}(\Gamma^{\mathrm{sep}})$. These expressions correspond to paths from $[1]$ to $[g]$ in $\mathrm{Wh}(\Gamma)$, and the union of these paths connects $\mathrm{Wh}(\Gamma)$. ∎
Consider $\mathrm{Wh}(\mathbf{R}_{n})$. This space has a vertex for every conjugacy class of elements of $\mathbf{F}_{n},$ and its edges are labeled by conjugacy classes of nontrivial elements of $\mathcal{C}^{\mathrm{sep}}$. Since all primitive elements of $\mathbf{F}_{n}$ are in $\mathcal{C}^{\mathrm{sep}}$, the space $\mathrm{Wh}(\mathbf{R}_{n})$ is connected.
Theorem 3.4.
$\mathrm{Wh}(\mathbf{R}_{n})$ admits an $\mathrm{Out}(\mathbf{F}_{n})$-action by isometries.
Proof.
First we show $\mathrm{Wh}(\mathbf{R}_{n})$ admits an $\mathrm{Aut}(\mathbf{F}_{n})$-action. Let $\phi\in\mathrm{Aut}(\mathbf{F}_{n})$. The automorphism $\phi$ sends $[v]$ to $[\phi(v)]$. Suppose there is an edge $[\alpha]\in\mathrm{Wh}(\mathbf{R}_{n})$ from $[v]$ to $[w]$. Then there are elements $c,d\in\mathbf{F}_{n}$ satisfying $vc\alpha\overline{c}=dw\overline{d}$. If we apply $\phi$ to this equation, we get $\phi(v)\phi(c)\phi(\alpha)\overline{\phi(c)}=\phi(d)w\overline{\phi(d)}.$ So in $\mathrm{Wh}(\mathbf{R}_{n})$, the image of the edge $[\alpha]$ with endpoints $[v],[w]$ is an edge $[\phi(\alpha)]$ with endpoints $[\phi(v)],[\phi(w)]$. Since $\phi\in\mathrm{Aut}(\mathbf{F}_{n})$ is an automorphism, it permutes conjugacy classes of separable words, hence $\phi(\alpha)$ is separable. Note that if $\phi$ is an inner automorphism, $\phi$ fixes $\mathrm{Wh}(\mathbf{R}_{n})$ pointwise, so we can upgrade the $\mathrm{Aut}(\mathbf{F}_{n})$-action on $\mathrm{Wh}(\mathbf{R}_{n})$ to an $\mathrm{Out}(\mathbf{F}_{n})$-action.
The action of $\mathrm{Out}(\mathbf{F}_{n})$ preserves the metric on $\mathrm{Wh}(\mathbf{R}_{n})$: let $\phi\in\mathrm{Out}(\mathbf{F}_{n})$ and ${[v],[w]\in\mathrm{Wh}(\mathbf{R}_{n})}$. Let $p([v],[w])$ be any path between $[v]$ and $[w]$. Since automorphisms of the free group preserve separability, $||v||_{\mathrm{sep}}=||\phi(v)||_{\mathrm{sep}}$, and $\mathrm{length}(p)=\mathrm{length}(\phi(p))$. The action of $\mathrm{Out}(\mathbf{F}_{n})$ preserves $1$, so the action of $\mathrm{Out}(\mathbf{F}_{n})$ on the Whitehead complex of the rose is an action on a rooted graph. ∎
Remark 3.5.
Note that the only vertex of $\mathrm{Wh}(\mathbf{R}_{n})$ fixed by $\mathrm{Out}(\mathbf{F}_{n})$ is $[1]$, so by deleting $[1]$ from $\mathrm{Wh}(\mathbf{R}_{n})$, we get a space with an isometric action of $\mathrm{Out}(\mathbf{F}_{n})$ with no fixed points. It is convenient for us to include $[1]$, but the proofs of all statements about $\mathrm{Wh}(\mathbf{R}_{n})$ could be altered to apply to $\mathrm{Wh}(\mathbf{R}_{n})\setminus\{[1]\}.$
We get a similar result to 3.4 for all $\mathrm{Wh}(\Gamma)$ where $\Gamma$ represents a characteristic subgroup of $\mathbf{F}_{n}$. We do not get an $\mathrm{Out}(\mathbf{F}_{n})$-action on the Whitehead complexes of graphs representing subgroups which are not characteristic: by definition, subgroups that are not characteristic are not preserved under the action of $\mathrm{Aut}(\mathbf{F}_{n})$ on $\mathbf{F}_{n}$. Since $\Gamma$ represents a normal subgroup of $\mathbf{F}_{n}$, $\pi_{1}(\Gamma)$ is fixed by $\mathrm{Inn}(\mathbf{F}_{n})$, so there is an automorphism $\phi\in\mathrm{Out}(\mathbf{F}_{n})$ and $g\in\pi_{1}(\Gamma)$ such that $\phi(g)\not\in\pi_{1}(\Gamma)$.
There is a connected component of $\mathrm{Wh}(\Gamma)$ containing all nontrivial separable conjugacy classes of $\pi_{1}(\Gamma)$. If $\Gamma$ represents a characteristic subgroup, this “separable component” is preserved under the action of $\mathrm{Out}(\mathbf{F}_{n})$.
Theorem 3.6.
Let $\Gamma$ represent a characteristic subgroup of $\mathbf{F}_{n}$. Then $\mathrm{Wh}(\Gamma)$ admits an isometric action of $\mathrm{Out}(\mathbf{F}_{n})$. This action fixes the component of $\mathrm{Wh}(\Gamma)$ containing all elements of $\pi_{1}(\Gamma^{\mathrm{sep}})$.
Proof.
For the first statement, note that since $\Gamma$ is characteristic, $\pi_{1}(\Gamma)$ is preserved under the action of $\mathrm{Out}(\mathbf{F}_{n})$ on $\mathbf{F}_{n}$. The rest of the proof is the same as the proof of 3.4.
For the second statement, let $C$ be the component of $\mathrm{Wh}(\mathbf{R}_{n})$ containing the trivial vertex. Then $C$ contains vertices for all $[g]\subset\pi_{1}(\Gamma^{\mathrm{sep}})$ as there are edges from $[1]$ to the conjugacy classes of every separable element of $\pi_{1}(\Gamma)$. Let $[y]\in C$, so there is a path ${p=[\alpha_{1}],[\alpha_{2}],\ldots,[\alpha_{m}]}$ from the trivial vertex to $[y]$ where $m\in\mathbb{N}$ and $\alpha_{i}\in\Gamma^{\mathrm{sep}}$. For ${\phi\in\mathrm{Out}(\mathbf{F}_{n})},$ $\phi(p)$ is a path with edges $[\phi(\alpha_{1})],[\phi(\alpha_{2})],\ldots,[\phi(\alpha_{m})]$. Since $\phi$ is an automorphism and $\Gamma$ is characteristic, $[\phi(\alpha_{i})]\in\Gamma^{\mathrm{sep}}$ for all $i,$ so $\phi(p)$ is a path from the trivial vertex to $[\phi(y)]$, implying ${[\phi(y)]\in C}$. ∎
Finally, regarding the topology of $\mathrm{Wh}(\mathbf{R}_{n})$, we have the following proposition.
Proposition 3.7.
The flag complex of $\mathrm{Wh}(\mathbf{R}_{n})$ is infinite-dimensional for all $n\geq 2$.
Proof.
Recall that we let $\mathbf{F}_{n}$ be generated by the set $X_{n}=\{x_{1},x_{2},\ldots,x_{n}\}$. Let ${Y=\{x_{1}x_{2}^{m}\}_{m=1}^{\infty}.}$ Every element of $Y$ is primitive in $\mathbf{F}_{n}$ for all $n$, so every element of $Y$ is separable. Without loss of generality, let $k$ and $l$ be positive integers such that $k<l$. Observe that $d_{\mathrm{sep}}\Big{(}[ab^{k}],[ab^{l}]\Big{)}=1$ for all $k,l$ as $ab^{k}\cdot b^{l-k}=ab^{l}.$ Thus, for all $d\geq 1$, $\mathrm{Wh}(\mathbf{R}_{n})$ contains the one-skeleton of a simplex of dimension $d$ on the vertices $\{[1],[ab],[ab^{2}],\ldots,[ab^{d}]\}.$
3.1. Diameter of the Whitehead complex
In this section, we prove that the diameter of every component of $\mathrm{Wh}(\Gamma)$ is infinite. As a corollary, we show that the diameters of $\mathrm{Cay}(\mathbf{F}_{n};\mathcal{C}^{\mathrm{sep}})$ and $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$ are infinite. The key idea of the proof is that paths in the Whitehead complex correspond to diagrams for conjugacy classes of $\mathbf{F}_{n}$. If $\mathrm{Wh}(\mathbf{R}_{n})$ had finite diameter, there would be a universal bound on the areas of minimal Van Kampen diagrams for every element of $\mathbf{F}_{n}$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$. In the proof, we utilize positive words.
Definition 3.8.
A word $w\in\mathcal{W}_{n}$ is positive if all of the letters in $w$ are positive.
We begin by proving two lemmas. First, we show that while a separable word can have inseparable subwords, these inseparable subwords are “barely” inseparable: in particular, inseparable subwords $v$ of separable words have the property that the graph resulting from removing a certain edge of $\Omega(v)$ has a cut vertex (see Lemma 3.9). In addition, we show that if $D$ is a Van Kampen diagram for a positive word, $D$ has no cut edges (see Lemma 3.11).
Lemma 3.9.
Let $y_{i}\in X_{n}$ and suppose $u=y_{1}y_{2}\ldots y_{l}$ is a positive word satisfying the property that removing the edge corresponding to $y_{l}y_{1}$ from the Whitehead graph of $u$ results in a graph with no cut vertices. Let $t$ be any positive word and $w=ut$. Then $w$ is inseparable.
Proof.
Let $t=z_{l+1}\ldots z_{m}$, where each $z_{i}$ is in $X_{n}$. The Whitehead graph of $w$ has all the edges in the Whitehead graph of $u$ except the edge corresponding to $y_{l}y_{1}$, plus the edges corresponding to $\{y_{l}z_{l+1},z_{j}z_{j+1},z_{m}y_{1}\mid l+1\leq j\leq m-1\}$. Adding edges to a graph with no cut vertices results in a graph with no cut vertices, so $w$ is inseparable. ∎
Definition 3.10.
A cut edge of a Van Kampen diagram $D$ is an edge $e\in D$ which satisfies the property that removing a single point on the interior of $e$ disconnects $D$.
Lemma 3.11.
Let $w$ be a positive word. Then if $D$ is a Van Kampen diagram for $w$, $D$ has no cut edges.
Proof.
If $D$ is a Van Kampen diagram for $w$ which contains a cut edge $e,$ then $w$ contains both $\mathrm{lab}(e)$ and $\mathrm{lab}(e)^{-1}$ as subwords, so $w$ is not positive. ∎
Let $\widehat{\mathcal{W}_{n}^{\infty}}$ be the set of freely reduced one-sided infinite words, i.e. the subset of $\prod_{i=1}^{\infty}X_{n}^{*}$ containing exactly the elements with no subwords $x_{i}\overline{x_{i}}$ for any $i$. Our next lemma shows that there are positive infinite words $W\in\widehat{\mathcal{W}_{n}^{\infty}}$ with the property that all subwords of $W$ of sufficient length are inseparable and not a subword of any separable word. Later, we will show that paths in $\mathrm{Wh}(\mathbf{R}_{n})$ corresponding to these infinite words have infinite diameter.
Lemma 3.12.
Let $k\in\mathbb{N}$ and $n\geq 2$. If $k>4n^{2}$, there are infinite words $W\in\widehat{\mathcal{W}_{n}^{\infty}}$ with the property that if $w_{i}$ is an initial subword of $W,$ and $u\subset w_{i}$ is a cyclic subword of $w_{i}$ of length at least $k$, then $u$ is not contained in any proper free factor of $\mathbf{F}_{n}$.
Proof.
Recall that we take $X_{n}=\{x_{i}\mid 1\leq i\leq n\}$ to be the generating set for $\mathbf{F}_{n}$. The edges in the Whitehead graph of $w\in\mathbf{F}_{n}$ are in bijection with length-two cyclic subwords of $w$. If $w$ is a positive word, then by construction of the Whitehead graph, the most complex Whitehead profile $w$ can have is the complete bipartite graph $B_{n}$, where the partition on the vertices of the profile is given by $\{x_{i}\in X_{n}\}\bigsqcup\{\overline{x_{i}}\mid x_{i}\in X_{n}\}$ (see Figure 10). Observe that for $n\geq 3$, $B_{n}$ has the property that removal of a single edge yields a graph with no cut vertices, so by Lemma 3.9, if $n\geq 3$, $B_{n}$ cannot be the Whitehead profile of any subword of a separable word.
Let $n\geq 3$. In $\mathbf{F}_{n}$, there are $n^{2}$ positive words $x_{i}x_{j}$ of length two. To construct $W$, let $k>4n^{2}$ and ensure that each of the $x_{i}x_{j}$ are present in every subword of $W$ of length at least $\frac{k}{2}$ (see (1) of Example 3.13 for an example of such a word for $n=4$). Let $w$ be an initial subword of $W$ with a cyclic subword $u$ of length at least $k$. If $u$ is a subword of $w$, we are done as the Whitehead profile of $u$ is $B_{n}$. If $u=ts$ where $t$ is a terminal subword of $w$ and $s$ is an initial subword of $w$, either $t$ or $s$ has length at least $\frac{k}{2}$, so we are done as the Whitehead profile of either $t$ or $s$ is $B_{n}$.
For $n=2$, removal of an edge from $B_{2}$ does yield a cut vertex, so we require that every length-$k$ subword of $W$ contains two copies of every positive length-two subword $x_{i}x_{j}$ (see (2) of Example 3.13 for an example). This way, if $w\subset W$ is an initial subword, $u$ is a cyclic subword of $w$ of length at least $k$, and $e\in\Omega(u)$ is an edge, $\sigma(\Omega(u)\setminus\{e\})$ has no cut vertices. ∎
Figure 10. Maximal Whitehead profiles of positive words for small $n$
Example 3.13.
  1. (1)
    Let $n=4$ and let $\{a,b,c,d\}$ be a basis for $\mathbf{F}_{4}$. Consider the infinite word $W:=\prod_{i=1}^{\infty}w,$ where $w=a^{2}b^{2}c^{2}d^{2}acbdcadba$. Then $W$ satisfies the conditions of Lemma 3.12.
  2. (2)
    Let $n=2$. Let $\{a,b\}$ be a basis for $\mathbf{F}_{2}$, $w=a^{2}b^{2}a$, and $W=\prod_{i=1}^{\infty}w$. Then $W$ satisfies the conditions of Lemma 3.12.
Definition 3.14.
A boundary arc of a diagram $D$ is a component of the intersection of the boundary of $D$ and the boundary of a single region $R\subset D$. Boundary arcs are labeled by subwords $u\subset\mathrm{lab}(\partial D)$ such that $u$ is also a subword of $\mathrm{lab}(\partial R)$. See Figure 11.
Figure 11. A diagram $D$. Every labeled arc on the boundary of $D$ is a boundary arc except the pink arcs $c$ and $g$. The arc $c$ is not a boundary arc because it does not lie on the boundary of a region (see Remark 2.8), and $g$ is not a boundary arc because it is not the intersection of the boundary of a region with the boundary of $D$.
By Remark 2.8, we glue regions along cycles labeled by cyclically reduced words, so cut edges of diagrams are not boundary arcs.
Let $\mathcal{P}=\langle X_{n}\mid R\rangle$ be a group presentation. Our next proposition gives an upper bound on the number of boundary arcs of a “nice” Van Kampen diagram over $\mathcal{P}$ of area $a$ (see below for the definition of nice). Let $(w_{i})_{i}$ be an infinite sequence of positive words of increasing length, and suppose that there is an upper bound on the areas of minimal diagrams $D_{i}$ for $w_{i}$ over $\mathcal{P}$. Then the maximal length of a boundary arc of $D_{i}$ approaches infinity as $i\rightarrow\infty.$
In the rest of the paper, it will be helpful to consider diagrams with no valence-two vertices. Of course, Van Kampen diagrams normally have many valence-two vertices, but a Van Kampen diagram with valence-two vertices is topologically equivalent to a 2-complex without them. Suppose $v$ is a valence-two vertex in a Van Kampen diagram $D$ with incident edges $e_{1}$ and $e_{2}$ satisfying $\tau(e_{1})=\iota(e_{2})=v$. We can eliminate $v$ by replacing the path of length two labeled by $e_{1}e_{2}$ with a single edge $e$ of length two labeled by $\mathrm{lab}(e_{1})\mathrm{lab}(e_{2}).$
Definition 3.15.
A diagram is nice if it has no valence-2 vertices. A nice Van Kampen diagram is a 2-complex resulting from eliminating valence-two vertices from a Van Kampen diagram in the manner described above.
Proposition 3.16.
Let $\mathcal{P}=\langle X\mid R\rangle$ be a presentation, $w\in\langle\langle R\rangle\rangle,$ and $D$ a nice Van Kampen diagram for $w$ over $\mathcal{P}$ of area $a\geq 2$. Then the number of boundary arcs of $D$ is at most $2(a-1)$.
Proof.
All diagrams in this proof are nice (see Definition 3.15). The proof is by induction on $a$. For $a=2,$ $D$ is homotopy equivalent to one of the diagrams in Figure 12, all of which have two boundary arcs.
Figure 12. Van Kampen diagrams of area two.
Suppose that, for every diagram of area $b$ where $b<a$, $D$ has at most $2(b-1)$ boundary arcs. Consider a diagram $D$ of area $a$. Either there is an edge contained in the boundary of two distinct regions, or there is not. If there is, removing this edge gives us a diagram with $a-1$ regions. Otherwise $D$ is a wedge of disks, each of which contains exactly one region, so removing one of these disks gives us a diagram with $a-1$ regions. Let $e$ be the edge removed from $D.$ Consider how many boundary arcs can be added to $D\setminus\{e\}$ by attaching $e$ to $D\setminus\{e\}$.
There are two cases: either $e$ is a circle or $e$ is a line segment. If $e$ is a line segment, it has two endpoints. These endpoints can be attached at existing vertices or by subdividing arcs to create new vertices. Attaching a line segment to two existing boundary vertices does not increase the number of boundary arcs of $D$: see (a), (b), and (c) of Figure 13. Subdividing one boundary arc once and attaching a line segment with one endpoint at a previously existing vertex and another endpoint at the new vertex increases the number of boundary arcs by at most one (see (d), (e), and (f) of Figure 13). Subdividing the boundary twice and attaching $e$ at the two new vertices increases the number of boundary arcs by at most two, one for each subdivision (see (g) and (h) of Figure 13).
If $e$ is a copy of $\mathbb{S}^{1}$, at most one boundary arc of $D\setminus\{e\}$ is subdivided when $e$ is attached, so the number of boundary arcs increases by at most two (one for the subdivision, one for $e$ itself). See diagrams (i), (j), (k), and (l) of Figure 13. ∎
Figure 13. In diagrams (a), (b), and (c), a dashed red arc is attached at two existing boundary vertices. In diagrams (d), (e), and (f), arcs are attached at one existing vertex and one new vertex. In diagrams (g) and (h), arcs are attached at two new vertices. In diagrams (i) and (j), a copy of $S^{1}$ is attached at an existing boundary vertex. Finally, in diagrams (k) and (l), a copy of $S^{1}$ is attached at a new vertex.
We now prove that diameter of $\mathrm{Wh}(\mathbf{R}_{n})$ is infinite. Consider the presentation ${\mathcal{P}=\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle}$ of the trivial group. Every expression for $w\in\mathbf{F}_{n}$ as a product of elements of $\mathcal{C}^{\mathrm{sep}}$ has an associated Van Kampen diagram $D$ for $w$ over $\mathcal{P}$. The key idea of the proof is that $\mathrm{Wh}(\mathbf{R}_{n})$ has finite diameter if and only if there is a universal bound on the areas of minimal diagrams for all words in $\mathbf{F}_{n}$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$. We then take a sequence of initial subwords $(w_{i})_{i}$ of infinite words given by Lemma 3.12 and consider a sequence of minimal nice Van Kampen diagrams $D_{i}$ for the $w_{i}$. We show that for large enough $i$, some boundary arc along a single region of $D_{i}$ has length at least $k$. This is a contradiction as subwords of $w_{i}$ of length at least $k$ are not subwords of any separable word.
See 1.2
Proof.
All diagrams in this proof are nice. First, we prove that $\mathrm{Wh}(\mathbf{R}_{n})$ has infinite diameter. Take $W\in\widehat{\mathcal{W}_{n}^{\infty}}$ satisfying the conditions of Lemma 3.12, so $W$ is positive, and for every initial subword $w\subset W$, every length-$k$ cyclic subword of $w$ is inseparable. Take a sequence $(w_{i})_{i=1}^{\infty}$ of initial subwords of $W$ of strictly increasing length, and suppose toward a contradiction that $\mathrm{Wh}(\mathbf{R}_{n})$ has finite diameter.
Let $m_{i}=||w_{i}||_{\mathrm{sep}}$ and $w_{i}=\prod_{j=1}^{m_{i}}\alpha_{i,j}$ be expressions for the $w_{i}$ as products of elements of $\mathcal{C}^{\mathrm{sep}}$. These products correspond to paths $p_{i}$ of length $||w_{i}||_{\mathrm{sep}}$ in $\mathrm{Wh}(\mathbf{R}_{n})$ with initial vertex $[1]$ and terminal vertex $[w_{i}]$. By the finite diameter assumption, there is a universal bound on the lengths of minimal paths between vertices, hence there is a universal bound $B$ on $||w_{i}||_{\mathrm{sep}}$. By Lemma 3.3, these expressions for $w_{i}$ correspond to nice diagrams $D_{i}$ for $w_{i}$, and the number of regions of $D_{i}$ is $||w_{i}||_{\mathrm{sep}}$. By 3.16, the number of boundary arcs of $D_{i}$ is bounded above by $2(B-1)$ for all $i$.
Since all $w_{i}$ are positive, $|w_{i}|=|\partial D_{i}|$. Choose $i$ large enough so that $|w_{i}|>2k(B-1)$. By the pigeonhole principle, there is some boundary arc of $D_{i}$ of length at least $k$, which contradicts the assumption that each region in this diagram has boundary label in $\mathcal{C}^{\mathrm{sep}}$. Thus no universal bound on the areas of the $D_{i}$ exists. This concludes the proof that $\mathrm{Wh}(\mathbf{R}_{n})$ has infinite diameter.
We move on to the proof for $\Gamma\neq\mathbf{R}_{n}$. Observe that every component $C$ of $\mathrm{Wh}(\Gamma)$ contains a vertex representing the conjugacy class of a positive word as $\Gamma$ represents a finite-index subgroup of $\mathbf{F}_{n}$. To see this, take a cyclically reduced $\hat{w}\in[w]\subset C$, and suppose it contains $\overline{y^{m}}$ as a subword, where $y\in X_{n}.$ Let $l$ be the minimal power of $y$ such that $y^{l}$ lifts to $\Gamma$. Replacing $\overline{y^{m}}$ in $w$ with $y^{l-m}$ reduces the number of negative letters of $w$. So, we can assume $w$ is a positive word such that $[w]\in C$. Let $v$ be a product of positive powers of elements of $X_{n}$ which lift to $\pi_{1}(\Gamma)$ such that the Whitehead graph of $v$ is connected, has no cut vertices, and has the property that removing any edge from the Whitehead graph of $v$ results in a Whitehead graph with no cut vertices. By Lemma 3.9, $v$ is not a subword of any separable word. Define $W=w\prod_{i=1}^{\infty}v$, let $k=2(|v|+|w|)$, and let $\{w_{i}\}_{i=1}^{\infty}$ be an infinite sequence of initial subwords of $W$ of strictly increasing length. Observe every subword of $W$ of length $\frac{k}{2}$ is not a subword of any separable word. Then, as in the proof that $\mathrm{Wh}(\mathbf{R}_{n})$ has infinite diameter, if there is some upper bound on the separable lengths of the $\{w_{i}\}$, then there is an $i$ large enough so that some boundary arc of a minimal Van Kampen diagram for $w_{i}$ over $\langle X_{n}\mid\pi_{1}(\Gamma^{\mathrm{sep}})\rangle$ has length at least $k$, a contradiction. ∎
A corollary of 1.2 is below.
See 1.3
Proof.
Observe that $\mathrm{Wh}(\mathbf{R}_{n})$ can be obtained as a quotient of $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$. First, identify all vertices $v,w\in\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ satisfying $\mathrm{lab}(v)\in[\mathrm{lab}(w)]$. Here, the label of ${v\in\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})}$ is the element of $\mathbf{F}_{n}$ it represents. The result is a graph with edges labeled by elements of $\mathcal{C}^{\mathrm{sep}}$ and vertices labeled by conjugacy classes of $\mathbf{F}_{n}$. Then, for all edges $e_{1},e_{2}$ in the resulting graph satisfying $\mathrm{lab}(e_{1})\in[\mathrm{lab}(e_{2})]$, $\iota(e_{1})=\iota(e_{2}),$ and $\tau(e_{1})=\tau(e_{2})$, identify $e_{1}$ and $e_{2}$. Relabel all edges $e$ in the resulting graph by $[\mathrm{lab}(e)]$ to obtain $\mathrm{Wh}(\mathbf{R}_{n})$. This quotient map is 1-Lipschitz, so since $\mathrm{Wh}(\mathbf{R}_{n})$ has infinite diameter, so does $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}}).$ The set $\mathcal{C}^{\mathrm{sep}}$ contains $\mathcal{C}^{\mathrm{prim}}$, so $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$ has infinite diameter, as well. ∎
4. Nonhyperbolicity of the Whitehead complex
In this section, we prove that $\mathrm{Wh}(\mathbf{R}_{n})$ is nonhyperbolic for $n>1$ by finding, for each $n$, an infinite set of arbitrarily fat geodesic triangles. Before diving into the technicalities we give a detailed outline of the proof to serve as a guide to the reader.
  1. Step 1:
    We find a particular infinite set of positive separable words $\{r_{i}(n)\}_{i=1}^{\infty}$ such that none of them is contained in a free factor of rank less than $n-1$. For $n>2$, we show that if $i\neq j$ and $r_{i}$ and $r_{j}$ are subwords of a cyclically reduced $w\in\mathbf{F}_{n}$, then $w$ is inseparable. For $n=2$, we show that if $j>i+1$ and $r_{i}$ and $r_{j}$ are subwords of a cyclically reduced $w\in\mathbf{F}_{n}$, then $w$ is inseparable (see Lemma 4.2).
  2. Step 2:
    For $n>2$, let
    $w(a,k,s,n):=\prod_{i=a}^{a+s-1}r_{i}(n)^{k},$
    and for $n=2$, let
    $w(a,k,s,n):=\prod_{i=0}^{s-1}r_{a+2i}(n)^{k}.$
    We show that, for all $n$, and large enough $a$ and $k$, the word $w(a,k,s,n)$ has separable length $s$ (see 4.3). To get the lower bound, let $D$ be a nice Van Kampen diagram for $w(a,k,s,n)$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$, and suppose toward a contradiction that $\mathrm{area}(D)<s$. By 3.16, Van Kampen diagrams of area less than $s$ have at most $2(s-2)$ boundary arcs. As a consequence of Step 1, we obtain an upper bound on the length of any boundary arc of $D$. Given sufficiently large $a$ and $k$, we show that the sum of the lengths of all of the boundary arcs of $D$ is less than the length of $w(a,k,s,n)$ to obtain a contradiction.
  3. Step 3:
    For $n>2$, let $b=a+s$, and for $n=2$, let $b=a+2s.$ We show that, for $a$ and $k$ large enough, the distance $d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s$ using a proof similar to the proof that $||w(a,k,s,n)||_{\mathrm{sep}}=s$. A path in $\mathrm{Wh}(\mathbf{R}_{n})$ from $[w(a,k,s,n)]$ to $[w(b,k,s,n)]$ corresponds to an annular diagram (see Definition 2.10) with boundaries labeled by $w(a,k,s,n)$ and $w(b,k,s,n)$. We suppose toward a contradiction that $A$ is an annular diagram of area less than $2s$ with boundaries labeled as above. As before, Step 1 gives an upper bound on the lengths of boundary arcs of $A$. Let $O$ be the total length of the intersection of the two boundaries of $A$. In 4.8, we show that there are $a$ and $k$ sufficiently large so that the total length of the boundary arcs of $A$ is less than $|w(a,k,s,n)|+|w(b,k,s,n)|-O$. This is a contradiction as the boundaries of $A$ are labeled by positive words, so the total length of all boundary arcs of $A$ is equal to $|w(a,k,s,n)|+|w(b,k,s,n)|-O$.
  4. Step 4:
[…]
    $[w(a,k,s,n)w(a+s,k,s,n)]=[w(a,k,2s,n)]$
    (see Figure 14). The latter path lies above a ball of radius $s$ around $1$ by 4.3. For $\delta<s$, the vertex $[w(a,k,2s,n)]$ lies outside the $\delta$-neighborhood of the sides $([1],[w(a,k,s,n)])$ and $([1],[w(b,k,s,n)])$. Since these fat triangles exist for all separable lengths, and since there are elements of $\mathbf{F}_{n}$ with arbitrarily large separable length, we conclude that there is no $\delta$ satisfying the $\delta$-thin condition for hyperbolicity, completing the proof of 1.4.
Figure 14. We will show that, for $n>2,$ this is a geodesic triangle on the vertices $\{[1],w(a,k,s,n),w(b,k,s,n)\}$ in $\mathrm{Wh}(\mathbf{R}_{n})$. The dashed line represents the boundary of a ball of radius $s$ centered at the vertex $[1]$. A similar triangle with slightly different vertex and edge labels exists for $n=2$. See Figure 20.
4.1. Step 1: Properties of $\mathbf{r_{i}(n)}$
For $n>3$ define:
$r_{i}(n):=(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2};$
for $n=3$, define:
$r_{i}(n):=\big{(}(x_{1}^{2}x_{2}^{i}x_{3})^{2}x_{2}^{2}\big{)}^{2};$
for $n=2,$ define:
$r_{i}(n):=(x_{1}^{i}x_{2})^{2}.$
We will refer to $r_{i}(n)$ as $r_{i}$ when $n$ is clear from context. The need for separate cases for $n=3$ and $n=2$ will become clear. Let $i\geq 0$ and $n>2$. The set
$\{x_{1},x_{2},\ldots,x_{n-1},x_{1}^{2}x_{2}^{2}\cdot\ldots\cdot x_{n-2}^{2}x_{n-1}^{i}x_{n}\}$
is a basis for $\mathbf{F}_{n}$. For $n=2$, $\{x_{1},x_{1}^{2}x_{2}\}$ is a basis for $\mathbf{F}_{n}$. Since $r_{i}$ is a product of squares of $n-1$ of these basis elements, $r_{i}$ is separable for all $n$. Note that we take $r_{i}$ to be a product of squares because this guarantees the presence of edges between $x_{j}$ and $\overline{x_{j}}$ for $j<n$ in the Whitehead profile of $r_{i}$. This will matter when we prove that no separable word can have both $r_{i}$ and $r_{j}$ as subwords if $i\neq j$ (see Lemma 4.2).
Let $\phi$ be the following automorphism of $\mathbf{F}_{n}$:
$\phi(x_{i})=\begin{cases}x_{n-1}x_{i}\overline{x_{n-1}}&\quad\text{if }i<n-1;\\ x_{n-1}&\quad\text{if }i=n-1;\\ x_{n}\overline{x_{n-1}}&\quad\text{if }i=n.\\ \end{cases}$
Recall that $\mathcal{W}_{n}$ is the set of finite, possibly unreduced words with letters in $X_{n}^{*}$ and that an automorphism $\sigma\in\mathrm{Aut}(\mathbf{F}_{n})$ induces a map $\tilde{\sigma}:\mathcal{W}_{n}\rightarrow\mathcal{W}_{n}$. For $w\in\mathcal{W}_{n}$, $\tilde{\sigma}(w)$ is the possibly freely unreduced image of $w$ under $\tilde{\sigma}$, and $\widehat{\tilde{\sigma}(w)}$ is the freely reduced representative of $\tilde{\sigma}(w)$ in $\mathcal{W}_{n}.$
Lemma 4.1.
Let $w\in\mathcal{W}_{n}$ be freely reduced and let $\tilde{\phi}$ be the map $\mathcal{W}_{n}\rightarrow\mathcal{W}_{n}$ corresponding to $\phi$. Then the only letters in $\tilde{\phi}(w)$ not present in $\widehat{\tilde{\phi}(w)}$ are copies of $x_{n-1}$ and its inverse.
Proof.
Suppose that $\tilde{\phi}(w)=w_{1}yv\overline{y}w_{2}$, where $y\in X_{n}^{*}\setminus\{x_{n-1},\overline{x_{n-1}}\}$ and $v,w_{1},w_{2}\in\mathcal{W}_{n}$. We will show that $y$ and $\overline{y}$ do not cancel each other upon free reduction of $\tilde{\phi}(w)$.
For all $j\leq n$, there is exactly one copy of $x_{j}$ in $\phi(x_{j})$. The only other letters in $\phi(x_{j})$ are, possibly, one copy of $x_{n-1}$ and one copy of $\overline{x_{n-1}}$. Thus if $y\neq x_{n-1}^{\pm 1}$ is a letter of $\tilde{\phi}(w)$, this letter is the image of a copy of $y$ in $w$. We claim that we can choose $y,\overline{y}$ such that $v$ is a (possibly unreduced) word in the letters $x_{n-1}^{\pm 1}$. Suppose there is some letter $z\neq x_{n-1}^{\pm 1}$ between the copies of $y$ and $\overline{y}$ in $\tilde{\phi}(w)$. Then either there is a pair $z,\overline{z}$ between $y$ and $\overline{y}$ in $\tilde{\phi}(w)$, or $y$ and $\overline{y}$ do not cancel each other. If there is such a $z,\overline{z}$, in order to show $y$ and $\overline{y}$ do not cancel each other, it suffices to show that no copies of $z,\overline{z}$ cancel each other. In this way, we can continue to find pairs of letters $z^{\prime},\overline{z^{\prime}}$ between $y$ and $\overline{y}$ until the only letters between some copy of $z^{\prime}$ and $\overline{z^{\prime}}$ in $\tilde{\phi}(w)$ are copies of $x_{n-1}^{\pm 1}.$ Thus we assume the only letters between $y$ and $\overline{y}$ in $\tilde{\phi}(w)$ are copies of $x_{n-1}^{\pm 1}$.
Let $j\neq{n-1}$. Then $\phi(x_{j})$ contains $x_{j}$, one copy of $\overline{x_{n-1}}$, and possibly one copy of $x_{n-1}$. The copy of $x_{n-1}$, if present, appears before $x_{j}$ in $\phi(x_{j})$, and the copy of $\overline{x_{n-1}}$ appears after $x_{j}$ in $\phi(x_{j})$. It follows that if $y,\overline{y}$ are letters of $\tilde{\phi}(w)$, $y,\overline{y}$ are letters of $w$. Since $w$ is freely reduced, there is at least one letter between $y$ and $\overline{y}$ in $w$. Since we chose $y,\overline{y}$ so that $v\subset\tilde{\phi}(w)$ consists only of the letters $x_{n-1}^{\pm 1}$, these are the only letters between $y$ and $\overline{y}$ in $w$, as well. Suppose $w$ contains the subword $yx_{n-1}^{p}\overline{y}$ where $p\in\mathbb{Z}$. Then $\tilde{\phi}(w)$ contains the subword $y\overline{x_{n-1}}x_{n-1}^{p}x_{n-1}\overline{y},$ which freely reduces to $yx_{n-1}^{p}\overline{y}$. Thus $y$ and $\overline{y}$ do not cancel each other. ∎
Our second lemma states that, for $n\geq 2$, no separable word can have subwords $r_{i}$ and $r_{j}$ if $i\neq j$ (or $j>i+1$ if $n=2$). It is used to bound the lengths of boundary arcs of diagrams over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$ with boundary $\prod_{i}^{m<\infty}r_{i}^{k}$. The bound is used in the proofs of 4.3 and 4.8 which, respectively, compute $||w(a,k,s,n)||_{\mathrm{sep}}$ and $d_{\mathrm{sep}}\Big{(}w(a,k,s,n),w(b,k,s,n)\Big{)}.$
Lemma 4.2.
Let $i,j$ satisfy $1<i<j$ and let $n>1$.
  1. (1)
    For $n\geq 3$ and $i\neq j$, no separable $w\in\mathbf{F}_{n}$ can have both $r_{i}$ and $r_{j}$ as disjoint subwords;
  2. (2)
    For $n=2$ and $j>i+1$, no separable $w\in\mathbf{F}_{n}$ can have both $r_{i}$ and $r_{j}$ as disjoint subwords.
Proof.
For all $n$, we will show that, even though the Whitehead profile of $w$ may contain a cut vertex, the Whitehead profile of the image of $w$ after $i$ steps of Whitehead’s algorithm does not. Therefore, $w$ is inseparable. We divide the proof into three cases: $n>3,n=3,$ and $n=2$.
  1. (1i)
    Let $n>3$ and assume $w\in\mathbf{F}_{n}$ has subwords $r_{i}$ and $r_{j}$ where $j>i$. A subgraph of the Whitehead profile of $w$ can be seen in Figure 15. Observe that the only cut vertex of this graph is $x_{n-1}.$ The partition of $X_{n}^{*}$ associated with this Whitehead profile is $\{x_{n-1},\overline{x_{n}}\}\bigsqcup\{x_{1},\overline{x_{1}},x_{2},\overline{x_{2}},\ldots,x_{n-2},\overline{x_{n-2}},\overline{x_{n-1}},x_{n}\}$, and the Whitehead automorphism associated with this partition is the $\phi$ defined above Lemma 4.1
    Figure 15. The Whitehead profiles of $r_{i}(n)$ for $n\geq 2$.
    We have:
    $\phi(r_{i})=x_{n-1}(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i-1}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2}\overline{x_{n-1}}.$
    Since $i>1$, the Whitehead profile of $\phi(r_{i})$ contains the Whitehead profile in Figure 15. A calculation shows that
    $\phi^{m}(r_{i})=x_{n-1}^{m}(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i-m}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2}\overline{x_{n-1}}^{m}.$
    The first step of Whitehead’s algorithm applies the automorphism $\phi$ to $w$. Note that, since $x_{n-1}$ is the only possible cut vertex of $\phi(w)$, Whitehead’s algorithm repeatedly applies the automorphism $\phi$ to $w$. If for some $m<i$, the Whitehead profile of $\phi^{m}(w)$ has no cut vertices, $w$ is inseparable.
    Now consider $\phi^{i}(w)$. Let
    $u_{m}(i):=(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{m-i}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2}.$
    If $\phi^{i}(w)$ contains $u_{i}(i)$ and $u_{j}(i)$ as subwords, $w$ is inseparable, as the Whitehead graph of $u_{i}(i)$ contains the Whitehead profile in Figure 16, the Whitehead graph of $u_{j}(i)$ contains the Whitehead profile in Figure 15, and the union of these Whitehead profiles is connected with no cut vertices. Note that because $w$ is reduced, $\overline{x_{1}}$ does not directly precede $r_{i}$ or $r_{j}$ in $w$; also, $\overline{x_{2}}$ does not directly follow $r_{i}$ or $r_{j}$ in $w$. By Lemma 4.1, no letters are canceled in $\phi^{i}(w)$ but $x_{n-1}^{\pm 1}$, so since $n-1>2$, $\phi^{i}(w)$ contains $u_{i}(i)$ and $u_{j}(i)$ as disjoint subwords.
    Figure 16. The Whitehead profiles of $u_{i}(i)$ for $n\geq 2.$
  2. (1ii)
    Now assume $n=3$. We have
    $\phi(r_{i})=\big{(}x_{2}(x_{1}^{2}x_{2}^{i-1}x_{3})^{2}x_{2}\big{)}^{2}.$
    A calculation shows that, for $m\leq i$,
    $\phi^{m}(r_{i})=\big{(}x_{2}^{m}(x_{1}^{2}x_{2}^{i-m}x_{3})^{2}x_{2}^{2-m}\big{)}^{2}.$
    Similar to the proof of (1i), we want to show that if $r_{i}$ and $r_{j}$ are disjoint subwords of $w$, then $\phi^{m}(w)$ contains disjoint subwords $u_{i}(m),u_{j}(m)$ satisfying the property that the union of the Whitehead profiles of $u_{i}(i)$ and $u_{j}(i)$ is connected with no cut vertices. Let
    $u_{m}(i)=(x_{1}^{2}x_{2}^{m-i}x_{3}^{2})^{2}x_{2}^{2}(x_{1}^{2}x_{2}^{m-i}x_{3})^{2}.$
    The Whitehead profiles of $u_{i}(i)$ and $u_{j}(i)$ are shown in Figure 16 and Figure 15, respectively. Note that their union is connected without cut vertices. By Lemma 4.1, the only letters that can be cancelled under powers of $\phi$ are $x_{n-1}$ and its inverse, so it suffices to show that the copy of $x_{2}^{2}$ is preserved by $\phi$. Note that
    $\phi(x_{3}x_{2}^{2}x_{1})=x_{3}\overline{x_{2}}x_{2}^{2}x_{2}x_{1}\overline{x_{2}},$
    and the conclusion follows.
  3. (2)
    For $n=2$, we have:
    $\phi(r_{i})=x_{1}(x_{1}^{i-1}x_{2})^{2}\overline{x_{1}}.$
    A calculation shows that, for $m\leq i,$
    $\phi^{m}(r_{i})=x_{1}^{m}(x_{1}^{i-m}x_{2})^{2}\overline{x_{1}}^{m}.$
    Again, we want to show that $\phi^{i}(w)$ contains $u_{i}(i)$ and $u_{j}(i)$, where
    $u_{m}(i)=x_{2}x_{1}^{m-i}x_{2}.$
    The Whitehead profiles of $u_{i}(i)$ and $u_{j}(i)$ are visible in Figure 16 and Figure 15, respectively. Again, their union is connected without cut vertices. Lemma 4.1 implies that the only letters cancelled by $\phi$ are $x_{1}$ and its inverse. Thus $u_{j}(i)$ and $u_{i}(i)$ are both subwords of $\phi^{i}(w)$, and the conclusion follows. ∎
4.2. Step 2: Separable length of $\mathbf{w(a,k,s,n)}$
Recall that for $n>3,$
$r_{i}(n):=(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2};$
for $n=3$,
$r_{i}(n):=\big{(}(x_{1}^{2}x_{2}^{i}x_{3})^{2}x_{2}^{2}\big{)}^{2};$
for $n=2,$
$r_{i}(n):=(x_{1}^{i}x_{2})^{2}.$
For all $n\geq 3$, we let
(1) $w(a,k,s,n):=\displaystyle\prod_{i=a}^{a+s-1}r_{i}^{k};$
for $n=2$, let
(2) $w(a,k,s,n):=\displaystyle\prod_{i=0}^{s-1}r_{a+2i}^{k}.$
When $n$ is clear from context, we will drop it from the notation. Recall that a boundary arc of $D$ is a connected component of the intersection of the boundary of a region of $D$ with the boundary of $D$. Sometimes we will refer to a “boundary arc of $R$”, where $R$ is a region of $D$: we mean a boundary arc of $D$ contained in $R$. For convenience, we will often refer to a subword of the label of a boundary arc as a subword of a boundary arc, conflating a boundary arc with its label.
Proposition 4.3.
Let $n\geq 2$. For sufficiently large $a=a(s)$ and $k=k(a,n,s)$, the separable length of $w(a,k,s,n)$ is $s$.
Proof.
We have $||w(a,k,1,n)||_{\mathrm{sep}}=1$ as $w(a,k,1,n)$ is separable for all $a,k$ and $n$, and, by Lemma 4.2, $w(a,k,2,n)$ is inseparable for all $a,k$ and $n$, so $||w(a,k,2,n)||_{\mathrm{sep}}\geq 2$. For $n\geq 3$, let $b=a+s$, and for $n=2$, let $b=a+2s$. If $n\geq 3$, then, for all $s$, there is a path in $\mathrm{Wh}(\mathbf{R}_{n})$ from $[1]$ to $[w(a,k,s,n)]$ of length $s$ with edge labels $[r_{i}^{k}]$, $a\leq i\leq b-1$ (see Equation 1). Similarly, if $n=2$, there is a path from $[1]$ to $[w(a,k,s,n)]$ of length $s$ with edge labels $[r_{i}^{k}]$, where $i$ ranges over the set $\{a,a+2,a+4,\ldots,b-2\}$ (see Equation 2). Thus, for all $a,k,s,n$, $||w(a,k,s,n)||_{\mathrm{sep}}\leq s$. We conclude $||w(a,k,2,n)||_{\mathrm{sep}}=2$ for all $n.$
The rest of the proof of the proposition is broken into three cases ($n>3,$ $n=3,$ and $n=2$). Suppose toward a contradiction that $||w(a,k,s,n)||_{\mathrm{sep}}\leq s-1.$ By 3.16, a Van Kampen diagram of area at most $s-1$ has at most $2(s-2)$ boundary arcs. Let $D$ be a Van Kampen diagram for $w(a,k,s,n)$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$ of area at most $s-1$. For $n>2,$ by Lemma 4.2, each region of $D$ has at most one of the elements of $\{r_{i}\}_{i=a}^{a+s-1}$ as a subword of its boundary relation. Likewise, for $n=2$, each region of $D$ has at most one of the elements of $\{r_{a+2i}\}_{i=0}^{s-1}$ as a subword of its boundary relation. Since $D$ is a Van Kampen diagram for a positive word, the sum of the lengths of boundary arcs of $D$ is equal to $|w(a,k,s,n)|$ (see Lemma 3.11). The contradiction we will obtain is that the sum of the lengths of all boundary arcs of $D$ is less than $|w(a,k,s,n)|$ for large enough values of $a$ and $k$.
We will complete the proof of the proposition assuming the following two claims, the first of which gives an upper bound for the length of any boundary arc of $D$ and the second of which gives an upper bound for the sum of the lengths of boundary arcs of $D$. We then prove the claims.
Claim 4.4.
Let $n\geq 2$, let $D$ be a either a Van Kampen diagram or an annular diagram over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$, and suppose $D$ has a boundary cycle $B$ labeled by $w(a,k,s,n)$. Let
  1. (1)
    $M_{n}(a):=(k+2)(6n+2a+2s)$ for $n>3;$
  2. (2)
    $M_{n}(a):=(k+2)(12+4a+4s)$ for $n=3$;
  3. (3)
    $M_{n}(a):=(k+2)(2a+4s)$ for $n=2$.
Then, for sufficiently large $a$ and $k$, every boundary arc of $B$ has length less than $M_{n}(a).$
Claim 4.5.
Let $n\geq 2$. Let $D$ be a Van Kampen diagram for $w(a,k,s,n)$ over the presentation $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$. Suppose $D$ has at most $s-1$ regions. If $a$ and $k$ are sufficiently large, the sum of the lengths of the boundary arcs of $D$ is less than $(s-1)M_{n}(a)+(s-3)m_{n}(a),$ where $m_{n}(a)$ is given below.
  1. (1)
    For $n>3,$ $m_{n}(a):=12n+4a+4s;$
  2. (2)
    For $n=3,$ $m_{n}(a):=18+6a+6s;$
  3. (3)
    For $n=2$, $m_{n}(a):=3a+6s.$
We now begin the proof assuming the claims. By 4.5, the sum of the lengths of the boundary arcs of $D$ is less than $(s-1)M_{n}(a)+(s-3)m_{n}(a).$ The proof splits into three cases: $n>3$, $n=3$, and $n=2$.
Case 1: $\mathbf{n>3}$. A calculation shows
$\begin{split}|w(a,k,s,n)|&=\displaystyle\sum_{i=a}^{a+s-1}\big{(}k|r_{i}|\big{)}\\ &=k(6ns-11s+2as+s^{2}).\end{split}$
By the definitions of $M_{n}(a)$ and $m_{n}(a),$ we have
$(s-1)M_{n}(a)+(s-3)m_{n}(a)=(s-1)(k+2)(6n+2a+2s)+(s-3)(12n+4a+4s).$
So, $(s-1)M_{n}(a)+(s-3)m_{n}(a)<|w(a,k,s,n)|$ if and only if
(3) $(s-1)(k+2)(6n+2a+2s)+(s-3)(12n+4a+4s)<k(6ns-11s+2as+s^{2})$
which happens if and only if
(4) $\begin{split}2(s-1)&(6n+2a+2s)+(s-3)(12n+4a+4s)\\ &<k\Big{[}6ns-11s+2as+s^{2}-(s-1)(6n+2a+2s)\Big{]}\end{split}$
The right side of Equation 4 can be simplified to
$k(6n+2a-9s-s^{2}).$
For $a$ satisfying $a>s^{2}+9s$, $6n+2a-9s-s^{2}$ is positive. To satisfy Equation 3, take $k$ such that
(5) $k>\frac{(4s-8)(6n+2a+2s)}{6n+2a-9s-s^{2}}.$
Thus, choosing $k$ so that Equation 5 holds, we arrive at the desired contradiction.
Case 2: $\mathbf{n=3}$. Recall that
$r_{i}(3)=\big{(}(x_{1}^{2}x_{2}^{i}x_{3})^{2}x_{2}^{2}\big{)}^{2}$
and
$w(a,k,s,3)=\displaystyle\prod_{i=a}^{a+s-1}r_{i}(3)^{k}.$
Observe $|r_{i}(3)|=16+4i$. A calculation shows that
$\begin{split}|w(a,k,s,3)|&=\displaystyle\sum_{i=a}^{a+s-1}|r_{i}^{k}|\\ &=k(14s+4as+2s^{2}).\end{split}$
By the definitions of $M_{3}(a)$ and $m_{3}(a),$ we then have
$(s-1)M_{3}(a)+(s-3)m_{3}(a)=(s-1)(k+2)(12+4a+4s)+(s-3)(18+6a+6s).$
Now, $(s-1)M_{3}(a)+(s-3)m_{3}(a)<|w(a,k,s,3)|$ if and only if
(6) $k(14s+4as+2s^{2})>(s-1)(k+2)(12+4a+4s)+(s-3)(18+6a+6s),$
which happens if and only if
(7) $k(14s+4as+2s^{2})-k(s-1)(12+4a+4s)>2(s-1)(12+4a+4s)+(s-3)(18+6a+6s).$
The left side of Equation 6 simplifies to $2k(3s-s^{2}+6+2a),$ which is positive if $a>s^{2}.$ Thus, letting $a>s^{2}$ and
(8) $k>\displaystyle\frac{(7s-13)(6+2a+2s)}{2(3s-s^{2}+6+2a)}$
implies Equation 6 holds, which in turn implies the area of a Van Kampen diagram for $w(a,k,s,3)$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$ is at least $s.$
Case 3: $\mathbf{n=2}$. Recall that
$w(a,k,s,2)=\displaystyle\prod_{j=0}^{s-1}r_{a+2j}(2)^{k}$
and
$r_{i}(2)=(x_{1}^{i}x_{2})^{2}.$
The length of $r_{i}(2)$ is $22i+$ 2, so, after a calculation, we obtain
$\begin{split}|w(a,k,s,2)|&=k\displaystyle\sum_{j=0}^{s-1}\big{(}2(a+2j)+2\big{)}\\ &=k(2as+2s^{2}).\end{split}$
By definition of $M_{2}(a)$ and $m_{2}(a)$,
$(s-1)M_{2}(a)+(s-3)m_{2}(a)=(s-1)(k+2)(2a+4s)+(s-3)(3a+6s).$
As before, it suffices to show there are $a$ and $k$ large enough so that
$|w(a,k,s,2)|>(s-1)M_{2}(a)+(s-3)m_{2}(a).$
This is true if and only if
(9) $k(2as+2s^{2})>(s-1)(k+2)(2a+4s)+(s-3)(3a+6s),$
which is true if and only if
(10) $k(2as+2s^{2})-k(s-1)(2a+4s)>2(s-1)(2a+4s)+(s-3)(3a+6s).$
The left side of Equation 10 simplifies to $k(4s+2a-2s^{2})$, which is positive if $a>s^{2}.$ Letting $a>s^{2}$ and
(11) $k>\displaystyle\frac{(7s-13)(a+2s)}{4s+2a-2s^{2}}$
ensures that Equation 9 holds. This completes the proof of 4.3 assuming 4.4 and 4.5. We now prove the claims.
See 4.4
Proof.
Case 1: $\mathbf{n>3}$. Recall that
$r_{i}(n)=(x_{1}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}\ldots x_{2}^{2}$
and
$w(a,k,s,n)=\displaystyle\prod_{i=a}^{a+s-1}r_{i}(n)^{k}.$
The length of each boundary arc of $B$ is bounded above by the length of the longest subword of $w(a,k,s,n)$ containing only one $r_{i}$ as a subword, which is less than
$\max_{a\leq i\leq a+s-3}\{|r_{i}r_{i+1}^{k}r_{i+2}|,|r_{a+s-1}r_{a}^{k}r_{a+1}|,|r_{a+s-2}r_{a+s-1}^{k}r_{a}|\}.$
Since
$|r_{i}|$ $=|(x_{1}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}\ldots x_{2}^{2}|$
$=6n+2i-10,$
the word $r_{i}r_{i+1}^{k}r_{i+2}$ is shorter than $r_{j}r_{j+1}^{k}r_{j+2}$ when $i<j$. Therefore, the longest boundary arc of $B$ has length less than
$\max\{|r_{a+s-2}r_{a+s-1}^{k}r_{a}|,|r_{a+s-3}r_{a+s-2}^{k}r_{a+s-1}|,|r_{a+s-1}r_{a}^{k}r_{a+1}|\}.$
A calculation shows that the lengths of the three possible maxima are as follows:
$|r_{a+s-3}r_{a+s-2}^{k}r_{a+s-1}|$ $=|r_{a+s-3}|+k|r_{a+s-2}|+|r_{a+s-1}|$
$=12n+4a+4s-28+k(6n+2a+2s-14);$
$|r_{a+s-2}r_{a+s-1}^{k}r_{a}|$ $=|r_{a+s-2}|+k|r_{a+s-1}|+|r_{a}|$
$=12n+4a+2s-24+k(6n+2a+2s-12);$
$|r_{a+s-1}r_{a}^{k}r_{a+1}|$ $=|r_{a+s-1}|+k|r_{a}|+|r_{a+1}|$
$=12n+4a+2s-20+k(6n+2a-10).$
We can assume that $s>2$ by the first paragraph of the proof of this proposition. Now take $k$ such that $k>s-2$, and recall $s-2\geq 1$. This guarantees that $|r_{a+s-2}r_{a+s-1}^{k}r_{a}|$ is longest. To see this, consider our three possible maxima:
  1. (1)
    $|r_{a+s-3}r_{a+s-2}^{k}r_{a+s-1}|=12n+4a+4s-28+k(6n+2a+2s-14);$
  2. (2)
    $|r_{a+s-2}r_{a+s-1}^{k}r_{a}|=12n+4a+2s-24+k(6n+2a+2s-12)$;
  3. (3)
    $|r_{a+s-1}r_{a}^{k}r_{a+1}|=12n+4a+2s-20+k(6n+2a-10).$
Subtracting the third expression from each of the possible maxima, we obtain:
  1. (a)
    $2s-8+k(2s-4);$
  2. (b)
    $-4+k(2s-2);$
  3. (c)
    00.
Since $s>2$ and $k>s-2\geq 1$, both $(a)$ and $(b)$ above are positive. We also have
$2s-8+k(2s-4)$ $<-4+k(2s-2)$
$\iff 2s-8-2ks-4k$ $<-4+2ks-2k$
$\iff 2s-4$ $<2k$
$\iff s-2$ $<k.$
Thus, for $s>2$, $k>s-2$, and $n>3$, the longest boundary arc of $B$ has length less than
$12n+4a+2s-24+k(6n+2a+2s-12).$
Note
$12n+4a+2s-24+k(6n+2a+2s-12)<(k+2)(6n+2a+2s)=:M_{n}(a).$
Case 2: $\mathbf{n=3}$. Recall that
$r_{i}(3)=\big{(}(x_{1}^{2}x_{2}^{i}x_{n})^{2}x_{2}^{2}\big{)}^{2}$
and
$w(a,k,s,3)=\displaystyle\prod_{i=a}^{a+s-1}r_{i}(3)^{k}.$
As in Case 1, the length of each boundary arc of $B$ is bounded above by the length of the longest subword of $w(a,k,s,n)$ containing only one $r_{i}$ as a subword, which is less than
$\displaystyle\max_{a\leq i\leq a+s-3}\{|r_{i}r_{i+1}^{k}r_{i+2}|,|r_{a+s-2}r_{a+s-1}^{k}r_{a}|,|r_{a+s-1}r_{a}^{k}r_{a+1}|\}.$
Since
$|r_{i}(3)|=|\big{(}(x_{1}x_{2}^{i}x_{3})^{2}x_{2}^{2}\big{)}^{2}|=16+4i,$
the word $r_{i}r_{i+1}^{k}r_{i+2}$ is shorter than $r_{j}r_{j+1}^{k}r_{j+2}$ when $i<j$, so the longest boundary arc of $D$ along $w(a,k,s,n)$ has length less than
$\max\{|r_{a+s-2}r_{a+s-1}^{k}r_{a}|,|r_{a+s-3}r_{a+s-2}^{k}r_{a+s-1}|,|r_{a+s-1}r_{a}^{k}r_{a+1}|\}.$
We compute the lengths of the three possible maxima:
$|r_{a+s-3}r_{a+s-2}^{k}r_{a+s-1}|$ $=|r_{a+s-3}|+k|r_{a+s-2}|+|r_{a+s-1}|$
$=16+8a+8s+k(8+4a+4s);$
$|r_{a+s-2}r_{a+s-1}^{k}r_{a}|$ $=|r_{a+s-2}|+k|r_{a+s-1}|+|r_{a}|$
$=24+8a+4s+k(12+4a+4s);$
$|r_{a+s-1}r_{a}^{k}r_{a+1}|$ $=|r_{a+s-1}|+k|r_{a}|+|r_{a+1}|$
$=32+8a+4s+k(16+4a).$
A calculation similar to that in Case 1 shows that, since $s>2$ and $k>s-2,$ the longest of these is $r_{a+s-2}r_{a+s-1}^{k}r_{a}.$ Thus, for $s>2$, $k>s-2$, and $n=3$, the longest boundary arc of $B$ has length less than
$24+8a+4s+k(12+4a+4s).$
Observe
$24+8a+4s+k(12+4a+4s)<(k+2)(12+4a+4s)=:M_{3}(a).$
Case 3: n=2. The proof mirrors the proofs of the other two cases. Recall that
$r_{i}(2)=(x_{1}^{i}x_{2})^{2}$
and
$w(a,k,s,2)=r_{a}^{k}r_{a+2}^{k}r_{a+4}^{k}\ldots r_{a+2(s-1)}.$
We calculate
$|r_{i}(2)|=|(x_{1}^{i}x_{2})^{2}|=2+2i.$
The longest subword of $w(a,k,s,2)$ with only one of the $\{r_{a+2j}\}_{j=0}^{s-1}$ as a subword has length less than
$\displaystyle\max_{a\leq i\leq a+2(s-3)}\{|r_{i}r_{i+2}^{k}r_{i+4}|,|r_{a+2(s-2)}r_{a+2(s-1)}^{k}r_{a}|,|r_{a+2(s-1)}r_{a}^{k}r_{a+2}|\}$
As before, $r_{i}r_{i+2}^{k}r_{i+4}$ is shorter than $r_{j}r_{j+2}r_{j+4}$ when $i<j$. Thus the longest boundary arc of $B$ has length less than
$\max\{|r_{a+2(s-3)}r_{a+2(s-2)}^{k}r_{a+2(s-1)}|,|r_{a+2(s-2)}r_{a+2(s-1)}^{k}r_{a}|,|r_{a+2(s-1)}r_{a}^{k}r_{a+2}|\}$
A calculation shows that, for for $s>2,k>s-2$, $|r_{a+2(s-2)}r_{a+2(s-1)}^{k}r_{a}|$ is largest. We conclude
$4a+4s-4+k(2a+4s-2)$
is larger than the length of the longest boundary arc of $B$. Note that
$4a+4s-4+k(2a+4s-2)<(k+2)(2a+4s)=:M_{2}(a).$
See 4.5
Proof of 4.5.
Recall that, for $n>3,$
$r_{i}(n)=(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2}$
and
$w(a,k,s,n)=\displaystyle\prod_{i=a}^{a+s-1}\big{(}(x_{1}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}\ldots x_{2}^{2}\big{)}^{k}.$
For $n=3,$
$r_{i}(n)=\big{(}(x_{1}^{2}x_{2}^{i}x_{3})^{2}x_{2}^{2}\big{)}^{2}$
and
$w(a,k,s,n)=\displaystyle\prod_{i=a}^{a+s-1}\big{(}(x_{1}^{2}x_{2}^{i}x_{3})^{2}x_{2}^{2}\big{)}^{2k}.$
For $n=2,$
$r_{i}(2)=(x_{1}^{i}x_{2})^{2}$
and
$w(a,k,s,n)=\displaystyle\prod_{j=0}^{s-1}(x_{1}^{a+2j}x_{2})^{2k}.$
Observe that, for all $n>2$, the only instance of $r_{i}$ as a cyclic subword of $w(a,k,s,n)$ is in the term $r_{i}^{k}$. For $n=2$, this is almost true: for every $i\neq a$, the only instance of $r_{i}$ as a cyclic subword of $w(a,k,s,n)$ is in the term $r_{i}^{k}$, and the only instance of $r_{a}$ as a cyclic subword of $w(a,k,s,n)$ is within $x_{1}^{a+2(s-1)}x_{2}r_{a}^{k}$. Thus if a region $R$ of $D$ has a boundary arc containing $r_{i}$ as a subword, $R$ intersects this boundary within the section labeled $r_{i}^{k}$ (or within $x_{1}^{a+2(s-1)}x_{2}r_{a}^{k}$ if $n=2$ and $i=a$). By Lemma 4.2, each region of $D$ contains at most one $r_{i}$ as a subword of its boundary relation. 4.4 implies that every boundary arc of $D$ containing an $r_{i}$ as a subword has length less than $M_{n}(a)$. Hence if $R$ is a region of $D$ and $i$ is an integer, the total length of the boundary arcs of $R$ which contain a subword $r_{i}$ is less than $M_{n}(a)$. By assumption, $D$ has at most $s-1$ regions, so the total length of the boundary arcs of $D$ containing an $r_{i}$ as a subword is at most $(s-1)M_{n}(a).$ The remaining boundary arcs of $D$ do not contain a subword $r_{i}$ for any $i$. For $n>3,$ the longest subwords of $w(a,k,s,n)$ containing no $r_{i}$ are of length less than
$|r_{a+s-1}(n)^{2}|=12n+4a+4s-24<12n+4a+4s=:m_{n}(a).$
For $n=3$, the longest subwords of $w(a,k,s,n)$ with no $r_{i}(3)$ as a subword are of length less than
$m_{3}(a):=|\big{(}(x_{1}^{2}x_{2}^{a+s-1}x_{3})^{2}x_{2}^{2}\big{)}^{3}|=18+6a+6s.$
For $n=2,$ the longest subwords of $w(a,k,s,n)$ with no $r_{i}(2)$ as a subword are of length less than
$|(x_{1}^{a+2s-2}x_{2})^{3}|=3a+6s-3<3a+6s=:m_{2}(a).$
Since $D$ has at most $2(s-2)$ boundary arcs, assuming each of the $s-1$ regions of $D$ contains a boundary arc with an $r_{i}$ subword, there are at most $(s-3)$ boundary arcs of $D$ with no $r_{i}$ subword. Thus, the sum of the lengths of boundary arcs of $D$ is less than $(s-1)M_{n}(a)+(s-3)m_{n}(a)$. This completes the proof of 4.5. ∎
With the proofs of both 4.4 and 4.5 complete, we have completed the proof of 4.3.∎
4.3. Step 3: Computing $\mathbf{d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}}$.
For $n>2$, let $b=a+s$, and for $n=2$, let $b=a+2s.$ We have shown so far that if $a$ and $k$ are large enough, the words $w(a,k,s,n)$ and $w(b,k,s,n)$ have separable length $s$. Now we want to show that we can choose $a$ and $k$ large enough so that $d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s$. The following lemma gives a way to visualize paths between positive words in $\mathrm{Wh}(\mathbf{R}_{n}),$ similar to Lemma 3.3.
Lemma 4.6.
Let $[v]$ and $[w]$ have positive representatives $\hat{v}$ and $\hat{w}$, respectively. If there is a path $p([v],[w])\in\mathcal{P}\mathrm{Wh}(\Gamma)$ between $[v]$ and $[w]$, there is an annular diagram $A_{p}$ over $\langle X_{n}\mid\pi_{1}(\Gamma^{\mathrm{sep}})\rangle$ with boundaries $\hat{v}$ and $\hat{w}$. The regions of $A_{p}$ have boundaries labeled by representatives of the conjugacy classes labeling the edges of $p$.
Proof.
Let the edges of $p([v],[w])$ be $[\alpha_{1}],[\alpha_{2}],\ldots,[\alpha_{l}]$ where $l$ is the length of $p([v],[w]).$ By the definition of $\mathrm{Wh}(\Gamma)$, there are $v^{\prime}\in[v]$ and $\hat{\alpha_{i}}\in[\alpha_{i}]$ so that $v^{\prime}\hat{\alpha_{1}}\hat{\alpha_{2}}\ldots\hat{\alpha_{l}}=\hat{w}$. To construct an annular diagram with boundaries $\hat{v}$ and $\hat{w}$, wedge disks with the same orientation and boundary labels $v^{\prime},\hat{\alpha_{1}},\hat{\alpha_{2}},\ldots,\hat{\alpha_{l}}$ at a point to obtain a 2-complex $C$, then fold until each disk has completely folded boundary and the boundary of the 2-complex is labeled by $\hat{w}$. The result is a Van Kampen diagram for $\hat{w}$ over $\langle X_{n}\mid v^{\prime},\mathcal{C}^{\mathrm{sep}}\rangle$. Then delete any degree-one vertices and the open disk with boundary $v^{\prime}$ to obtain the desired annular diagram. See Figure 17 for an example of an annular diagram corresponding to a specific path in $\mathrm{Wh}(\mathbf{R}_{n}).$
Figure 17. A path from $[a,b]$ to $[abc^{2}\overline{a}c^{2}\overline{b}abcab\overline{c}]$ in $\mathrm{Wh}(\mathbf{R}_{n})$ and a corresponding annular diagram.
Our proof that there are $a$ and $k$ such that
$d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s$
is similar to the proof that there are $a$ and $k$ such that $||w(a,k,s,n)||_{\mathrm{sep}}=s.$ First, we find an upper bound on the number of boundary arcs of annular diagrams of area less than $2s$. Then we use Lemma 4.2 to show that if $a$ and $k$ are sufficiently large and $A$ is an annular diagram with boundaries $w(a,k,s,n)$ and $w(b,k,s,n)$, and the area of $A$ is less than $2s$, then the sum of the lengths of the boundary arcs of $A$ is strictly less than $|w(a,k,s,n)|+|w(b,k,s,)|-O_{n},$ where $O_{n}$ is the total length of boundary arcs shared by $w(a,k,s,n)$ and $w(b,k,s,n).$ This is a contradiction as $w(a,k,s,n)$ and $w(b,k,s,n)$ are both positive, so every edge on the boundary of $A$ is either contained within the boundary of a region or is shared by both boundaries.
Figure 18. The three annular diagrams of area one.
Lemma 4.7.
Let $A$ be an annular diagram of area $a.$ Then the number of boundary arcs of $A$ is at most $2a$.
Proof.
All diagrams in this proof are nice (Definition 3.15). Like the proof of 3.16, we induct on $a$. If $a=1$, $A$ is one of the diagrams in Figure 18; note that each diagram has at most two boundary arcs. Suppose the conclusion holds for all annular diagrams of area less than $a$. If there are two regions of $A$ sharing an edge, remove this edge to decrease the area of the diagram by one. Otherwise, remove a boundary arc to decrease the area by one. Consider attaching an open interval to an annular diagram $A^{\prime}$ of area $a-1$. The interval could be attached at existing vertices of $A^{\prime}$, or it could be attached at new vertices obtained by subdividing edges of $A^{\prime}$. If we add an arc with both endpoints on one or two existing vertices, the number of boundary arcs of $A^{\prime}$ increases by at most one (see diagrams (a), (b), (c), and (d) of Figure 19). If we subdivide a boundary arc one time and add a boundary arc with endpoints at the new vertex and an existing vertex, the number of boundary arcs of $A^{\prime}$ increases by at most one (see diagrams (e), (f), and (g) of Figure 19). If we subdivide a boundary arc twice or subdivide two distinct boundary arcs and attach the endpoints of the interval to the two new vertices, the number of boundary arcs of $A^{\prime}$ increases by at most two (see diagrams (h), (i), (j), and (k) of Figure 19). Attaching the endpoints of an interval to vertices on the interior of $A^{\prime}$ does not increase the number of boundary arcs at all.
Figure 19. Adding dashed arcs to annular diagrams. In diagrams (a), (b), (c), and (d), arcs are attached at existing boundary vertices. In diagrams (e), (f), and (g), arcs are attached at one existing vertex and one new boundary vertex. In diagrams (h), (i), (j), and (k), arcs are attached at two new boundary vertices.
Proposition 4.8.
For $n>2$, let $b=a+s;$ for $n=2$, let $b=a+2s.$ Then, for large enough integers $a$ and $k$, $d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s.$
Proof.
By Lemma 4.6, we can view a path $p$ in $\mathrm{Wh}(\mathbf{R}_{n})$ from $[w(a,k,s,n)]$ to $[w(b,k,s,n)]$ as an annular diagram $A$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$ with boundaries $C$ labeled by $w(a,k,s,n)$ and $B$ labeled by $w(b,k,s,n)$. The number of components of $C\cap B$ is bounded above by the length of $p$ as the area of $A$ is the length of $p$. The components of $C\cap B$ are labeled by subwords shared by $w(a,k,s,n)$ and $w(b,k,s,n).$ By 4.3, for sufficiently large $a$ and $k,$ both $w(a,k,s,n)$ and $w(b,k,s,n)$ have separable length $s,$ so there is a path of length $2s$ from $w(a,k,s,n)$ to $w(b,k,s,n)$ through the trivial vertex.
It remains to show that there are $a$ and $k$ such that
$d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}\geq 2s,$
so suppose toward a contradiction that
$d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}<2s.$
Then there is an annular diagram $A^{\prime}$ over $\langle X_{n}\mid\mathcal{C}^{\mathrm{sep}}\rangle$ with boundaries $w(a,k,s,n)$ and $w(b,k,s,n)$ of area no more than $12s-$ 1. By Lemma 4.7, such a diagram has at most $2(2s-1)$ boundary arcs. We obtain a contradiction by computing upper and lower bounds for the length of the boundary of $A^{\prime}$ and showing that the upper bound for the length of the boundary of $A^{\prime}$ is less than the lower bound. The rest of the proof is broken into three cases: $n>3,$ $n=3$, and $n=2$.
Case 1: $\mathbf{n>3}$. By Lemma 4.2, no region of can have boundary label containing both $r_{i}$ and $r_{j}$ as subwords if $i\neq j$. By 4.4, the longest subword of $w(b,k,s,n)$ containing only one of the $r_{i}$ as a subword is of length less than $M_{n}(b)=(k+2)(6n+2b+2s)$. Subwords of $w(a,k,s,n)$ containing only one $r_{i}$ as a subword are strictly shorter. By 4.5, the longest subword of $w(b,k,s,n)$ containing no $r_{i}$ subword is of length strictly less than $m_{n}(b)=12n+4b+4s$. Subwords of $w(a,k,s,n)$ containing no $r_{i}$ subwords are strictly shorter. By the argument in 4.5, supposing $A^{\prime}$ contains two or more boundary arcs with the subword $r_{j}$ for some particular $j$, the sum of the lengths of these boundary arcs is less than $M_{n}(b)$. Thus the sum of the lengths of boundary arcs of $A^{\prime}$ containing any $r_{i}$ as a subword is less than $(2s-1)M_{n}(b).$ Assuming each region of $A^{\prime}$ contains a boundary arc with an $r_{i}$ subword, there are at most $12s-$ 1 remaining boundary arcs of $A^{\prime}$ with no $r_{i}$ as a subword, and these arcs have length less than $m_{n}(b)$. Thus the length of the boundary of $A^{\prime}$ is at most $(2s-1)M_{n}(b)+(2s-1)m_{n}(b).$
We now compute a lower bound on the length of the boundary of $A^{\prime}$ in order to find our contradiction. Consider the total length of the boundary of an annular diagram of area less than $s$ with boundaries labeled by $w(a,k,s,n)$ and $w(b,k,s,n)$. Recall that these boundaries overlap in at most $12s-$ 1 distinct arcs, and let $O_{n}$ be the length of the longest subword shared by $w(a,k,s,n)$ and $w(b,k,s,n)$. Since $w(a,k,s,n)$ and $w(b,k,s,n)$ are both positive, every edge of the boundary of $A^{\prime}$ is either contained in the boundary of a region of $A^{\prime}$ or is a subword of both $w(a,k,s,n)$ and $w(b,k,s,n).$ Thus the length of the boundary of $A^{\prime}$ is at least $|w(a,k,s,n)|+|w(b,k,s,n|-(2s-1)O_{n}$.
We want to compute an upper bound for $O_{n}$. Consider the longest subword shared by
$w(a,k,s,n)=\displaystyle\prod_{i=a}^{a+s-1}\big{(}(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2}\big{)}^{k}$
and
$w(a+s,k,s,n)=\displaystyle\prod_{i=a+s}^{a+2s-1}\big{(}(x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{i}x_{n})^{2}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2}\big{)}^{k}$
By inspection,
$O_{n}=|x_{n-1}^{a+s-1}x_{n}x_{n-1}^{2}x_{n-2}^{2}\ldots x_{2}^{2}x_{1}^{2}x_{2}^{2}\ldots x_{n-2}^{2}x_{n-1}^{a+s-1}|=2a+2s+4n-9.$
Thus, to show that there are $a$ and $k$ such that
$d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s,$
it suffices to show that there are $a$ and $k$ such that
(12) $|w(a,k,s,n)|+|w(a+s,k,s,n)|-(2s-1)O_{n}>(2s-1)M_{n}(b)+(2s-1)m_{n}(b).$
Recall that
$|w(a,k,s,n)|=k(6ns-11s+2as+s^{2}).$
A computation shows that
$|w(a,k,s,n)|+|w(b,k,s,n)|=k(12ns-22s+4as+4s^{2}).$
Therefore, Equation 12 is equivalent to
(13) $\begin{split}k(12ns&-22s+4as+4s^{2})-(2s-1)(2a+2s+4n-9)\\ &>(2s-1)(k+2)(6n+2a+4s)+(2s-1)(12n+4a+8s),\end{split}$
which simplifies to
$k(6n+2a-18s-4s^{2})>(2s-1)(28n+10a+18s-9).$
The left side of this inequality is positive if $a>4s^{2}+18s$. Therefore, letting $a>4s^{2}+18s$ and
$k>\displaystyle\frac{(2s-1)(28n+10a+18s-9)}{6n+2a-4s^{2}-18s}$
gives us our contradiction. Thus, for $n>3$ and sufficiently large $a$ and $k$,
$d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s.$
Case 2: $\mathbf{n=3}$. The structure of the proof is the same as above. We assume toward a contradiction that $A^{\prime}$ is an annular diagram of area at most $12s-$ 1 with boundaries $w(a,k,s,3)$ and $w(b,k,s,3)$. Let $O_{3}$ be the length of the longest subword shared by $w(a,k,s,3)$ and $w(b,k,s,3)$. Recall from Case 1 that $|w(a,k,s,3)|+|w(b,k,s,3)|-(2s-1)O_{3}$ is a lower bound for the length of the boundary of $A^{\prime}$, and an upper bound for the length of the boundary of $A^{\prime}$ is given by $(2s-1)M_{3}(b)+(2s-1)m_{3}(b)$. The contradiction we find is that there are $a$ and $k$ large enough so that
(14) $|w(a,k,s,3)|+|w(a+s,k,s,3)|-(2s-1)O_{3}>(2s-1)M_{3}(b)+(2s-1)m_{3}(b).$
A computation shows that
$\begin{split}|w(a,k,s,3)|+|w(a+s,k,s,3)|&=k(14s+4as+2s^{2})+k(14s+4s(a+s)+2s^{2})\\ &=k(28s+8as+8s^{2}).\end{split}$
We want to compute the length of the longest subword shared by
$w(a,k,s,3)=\Big{(}(x_{1}^{2}x_{2}^{a}x_{3})^{2}x_{2}^{2}\Big{)}^{k}\Big{(}(x_{1}^{2}x_{2}^{a+1}x_{3})^{2}x_{2}^{2}\Big{)}^{k}\ldots\Big{(}(x_{1}^{2}x_{2}^{a+s-1}x_{3})^{2}x_{2}^{2}\Big{)}^{k}$
and
$w(a+s,k,s,3)=\Big{(}(x_{1}^{2}x_{2}^{a+s}x_{3})^{2}x_{2}^{2}\Big{)}^{2k}\Big{(}(x_{1}^{2}x_{2}^{a+s+1}x_{3})^{2}x_{2}^{2}\Big{)}^{2k}\ldots\Big{(}(x_{1}^{2}x_{2}^{a+2s-1}x_{3})^{2}x_{2}^{2}\Big{)}^{2k}.$
By inspection, the longest subword of both $w(a,k,s,n)$ and $w(a+s,k,s,n)$ has length
$|x_{2}^{a+s-1}x_{3}x_{2}^{2}x_{1}^{2}x_{2}^{a+s-1}|=2a+2s+3$
so $O_{n}=2a+2s+3$. Thus a lower bound for the length of the boundary of $A^{\prime}$ is
$|w(a,k,s,n)|+|w(b,k,s,n)|-(2s-1)O_{n}=k(28s+8as+8s^{2})-(2s-1)(2a+2s+3).$
We now compute an upper bound for the length of the boundary of $A^{\prime}$. By the argument in Case 1, an upper bound for the length of the boundary of $A^{\prime}$ is given by
$(2s-1)M_{3}(b)+(2s-1)m_{3}(b).$
In 4.4 and 4.5, respectively, we computed that, for $w(b,s,k,n),$
$M_{3}(b)=(k+2)(12+4b+4s)$
and
$m_{3}(b)=18+6b+6s.$
Thus an upper bound for the length of the boundary of $A^{\prime}$ is
$(2s-1)(k+2)(12+4a+8s)+(2s-1)(18+6a+12s).$
To obtain a contradiction, we show that Equation 14 holds for sufficiently large $a$ and $k$. Equation 14 is equivalent to
(15) $\begin{split}k(28s+8as+8s^{2})-&(2s-1)(2a+2s+3)\\ &>(2s-1)(k+2)(12+4a+8s)+(2s-1)(18+6a+12s),\end{split}$
which simplifies to
(16) $k(12s-8s^{2}+12+4a)>(2s-1)(45+16a+30s).$
When $a>2s^{2}$, $12s-8s^{2}+12+4a$ is positive. So, letting $a>2s^{2}$ and
$k>\frac{(2s-1)(45+16a+30s)}{12s-8s^{2}+12+4a}$
gives us our contradiction. We conclude $d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s.$
Case 3: $\mathbf{n=2}$. The proof is the same as before. It suffices to show
(17) $|w(a,k,s,2)|+|w(a+2s,k,s,2)|-(2s-1)O_{2}>(2s-1)M_{2}(b)+(2s-1)m_{2}(b).$
Recall from 4.4 and 4.5 that, for $w(b,k,s,2)$, $M_{2}(b)=(k+2)(2b+4s)$ and $m_{2}(b)=3b+6s.$ 4.3 shows that $|w(a,k,s,2)|=k(2as+2s^{2})$, so we compute
$|w(a,k,s,2)|+|w(b,k,s,2)|=k(4as+8s^{2}).$
We want to compute $O_{2}$. Recall that
$w(a,k,s,2)=(x_{1}^{a}x_{2})^{2k}(x_{1}^{a+2}x_{2})^{2k}(x_{1}^{a+4}x_{2})^{2k}\ldots(x_{1}^{a+2s-2}x_{2})^{2k}$
and
$w(a+2s,k,s,2)=(x_{1}^{a+2s}x_{2})^{2k}(x_{1}^{a+2s+2}x_{2})^{2k}(x_{1}^{a+2s+4}x_{2})^{2k}\ldots(x_{1}^{a+4s-2}x_{2})^{2k}.$
By inspection, the longest shared subword of $w(a,k,s,2)$ and $w(b,k,s,2)$ is $x_{1}^{a+2s-2}x_{2}x_{1}^{a+2s-2},$ which has length $O_{2}=2a+4s-3.$ Equation 17 is equivalent to
(18) $k(4as+8s^{2})-(2s-1)(2a+4s-3)>(2s-1)(k+2)(2a+8s)+(2s-1)(3a+12s).$
After some simplification of Equation 18, we obtain
(19) $k(2a+8s-8s^{2})>(2s-1)(9a+32s-3).$
For $a>4s^{2}$, the left side of Equation 19 is positive. Thus, taking $a>4s^{2}$ and
(20) $k>\displaystyle\frac{(2s-1)(9a+32s-3)}{2a+8s-8s^{2}},$
gives us the desired contradiction. We conclude there are $a$ and $k$ large enough so that $d_{\mathrm{sep}}\Big{(}[w(a,k,s,2)],[w(b,k,s,2)]\Big{)}=2s.$
4.4. Step 4: Fat triangles.
We have proven that, for every $s>0$ we have geodesic triangles on the vertices $[w(a,k,s,n)],[w(b,k,s,n)],$ and $[1]$ where
$d_{\mathrm{sep}}\Big{(}[1],[w(a,k,s,n)]\Big{)}=d_{\mathrm{sep}}\Big{(}[1],[w(b,k,s,n)]\Big{)}=s$
and $d_{\mathrm{sep}}\Big{(}[w(a,k,s,n)],[w(b,k,s,n)]\Big{)}=2s.$ We want to show that there is a geodesic between $[w(a,k,s,n)]$ and $[w(b,k,s,n)]$ which lies above an open ball of radius $s$ centered at $[1]$. We conclude that these triangles violate the $\delta$-thin condition for hyperbolicity, so $\mathrm{Wh}(\mathbf{R}_{n})$ is nonhyperbolic for all $n\geq 2$.
See 1.4
Proof.
Let $n,s>0$. For $n>2,$ let $b=a+s$, and for $n=2$, let $b=a+2s$. For large enough $a$ and $k$, there is a geodesic $p$ from $w(a,k,s,n)$ to $w(b,k,s,n)$ with vertices $w(c_{i},k,t_{i},n)$ where all the $c_{i}\geq a$ and $t_{i}\geq s$ (see Figure 14). By 4.3, all vertices on the interior of $p$ have separable length strictly greater than $s$. This triangle is pictured in Figure 14 for $n>2$. For $n=2$, see Figure 20.
Figure 20. Fat geodesic triangles in $\mathrm{Wh}(\mathbf{R}_{2}).$
To show that $\mathrm{Wh}(\mathbf{R}_{n})$ is nonhyperbolic, it suffices to show that there is no constant $\delta>0$ such that the infinite collection of geodesic triangles on vertices $\{[1],[w(a,k,s,n)],[w(b,k,s,n)]\}_{s=1}^{\infty}$ described above satisfies the $\delta-$thin definition of hyperbolicity. But for any $\delta<\infty$, there is some $s>\delta$ such that $d_{\mathrm{sep}}\Big{(}[v],[w(a,k,2s,n)\Big{)}\geq s$ for any vertex $[v]$ on the geodesic segments $([1],[w(a,k,s,n)])$ and $([1],[w(b,k,s,n)])$ in Figure 14 and Figure 20. The conclusion follows. ∎
The geodesic triangles in 1.4 pull back to geodesic triangles in $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$, so we obtain the following result.
Figure 21. Geodesic triangle in $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ for $n>2$. Here, for $t,v\in\mathbf{F}_{n},$ $t^{v}:=vt\overline{v}.$
See 1.5
Proof.
Recall from the proof of 1.3 that $\mathrm{Wh}(\mathbf{R}_{n})$ is a quotient of $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ obtained by first identifying vertices labeled by conjugate elements of $\mathbf{F}_{n},$ then identifying edges between vertices which have conjugate labels. Let $\mathcal{T}$ be the set of the geodesic triangles in $\mathrm{Wh}(\mathbf{R}_{n})$ on the vertices
$\{[1],w(a,k,s,n),w(b,k,s,n)\}$
as described in the proof of 1.4. Let $T\in\mathcal{T}$ and note that the preimage of $T$ in $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ contains a triangle $\widetilde{T}$ (see Figure 21). The quotient map from $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ to $\mathrm{Wh}(\mathbf{R}_{n})$ is $1$-Lipschitz, so $\widetilde{T}$ is a geodesic triangle in $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$. It follows that the path from $w(a,k,s,n)$ to $w(b,k,s,n)$ crossing through $w(a,k,s,n)w(b,k,s,n)$ lies outside the ball of radius $s$ centered at $1\in\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$. Thus, there is an infinite family of geodesic triangles in $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ violating the $\delta$-thin condition for nonhyperbolicity.
To prove that $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$ is nonhyperbolic, it suffices to show that $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}})$ is quasi-isometric to $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$. To see this, let $w\in\mathcal{C}^{\mathrm{sep}}$. There is a basis $\mathcal{B}$ for $\mathbf{F}_{n}$ such that $w$ can be expressed as a word in at most $(n-1)$ elements of $\mathcal{B}$. Let $\mathcal{B}^{*}:=\mathcal{B}\cup\overline{\mathcal{B}}$, where $\overline{\mathcal{B}}$ is the set of inverses of elements of $\mathcal{B}$. Let $A$ be the set of elements of $\mathcal{B}^{*}$ present in $w$. The set *$\mathcal{B}^{*}\setminus A$ contains an element $a$ and its inverse $\overline{a}$. Write $w$ as a product of two nontrivial words $w_{1},w_{2}\in\mathbf{F}\langle\mathcal{B}\rangle$. Then $w_{1}a$ and $\overline{a}w_{2}$ are both primitive, since they only contain one copy of $a$ or its inverse (see Remark 2.6). Therefore, every separable word is a product of two primitive elements of $\mathbf{F}_{n}$, and it follows that $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{sep}})$ is quasi-isometric to $\mathrm{Cay}(\mathbf{F}_{n},\mathcal{C}^{\mathrm{prim}}),$ as desired. ∎
5. Homology version of the Whitehead complex
We define a nested sequence of spaces
$\mathrm{W}_{1}(\Gamma)\supset\mathrm{W}_{2}(\Gamma)\supset\mathrm{W}_{3}(\Gamma)\supset\ldots$
with intersection $\mathrm{Wh}(\Gamma)$. If $\Gamma$ is a regular cover of $\mathbf{R}_{n}$ and the deck group $G_{\Gamma}$ is nilpotent, this sequence will terminate after finitely many steps.
Definition 5.1.
Let $\Gamma$ be a finite cover of $\mathbf{R}_{n}$. We define a sequence of spaces of curves whose intersection is $\mathrm{Wh}(\Gamma)$. For all $i\geq 1$, we let $W_{i}(\Gamma)$ be the directed graph defined below.
  • The set of vertices of $W_{i}(\Gamma)$ is $\mathrm{Wh}(\Gamma)^{(0)};$
  • Edges are labeled by elements of the set $E_{i}:=\Gamma^{\mathrm{sep}}\cup\big{(}\pi_{1}(\Gamma)\big{)}_{i+1}$, where $\big{(}\pi_{1}(\Gamma)\big{)}_{j}\text{ is the }j$th term in the lower central series of $\pi_{1}(\Gamma)$;
  • There is an edge $[\alpha]\in E_{i}$ from $[v]$ to $[w]$ if and only if there are some $g\in[v],{h\in[w]},{c\in\mathbf{F}_{n}}$ such that $gc\alpha\overline{c}=h.$
Note that if any of the $W_{i}(\Gamma)$ are connected, so is $W_{1}(\Gamma)$. $W_{1}(\Gamma)$ is the “integral homology version” of the Whitehead complex. It is connected if and only if $\mathrm{H}_{1}^{\mathrm{sep}}(\Gamma;\mathbb{Z})=\mathrm{H}_{1}(\Gamma;\mathbb{Z}).$
See 1.6
Proof.
Suppose $\mathrm{H}^{\text{sep}}_{1}(\Gamma;\mathbb{Z})=\mathrm{H}_{1}(\Gamma;\mathbb{Z})$, and take an element $v$ of $\pi_{1}(\Gamma)$. Since $\mathrm{H}_{1}(\Gamma;\mathbb{Z})$ is the abelianization of $\pi_{1}(\Gamma)$, $v$ is expressible as a product $\prod_{i\leq m}\alpha_{i}$ such that ${\alpha_{i}\in\Gamma^{\mathrm{sep}}\cup[\pi_{1}(\Gamma),\pi_{1}(\Gamma)]}.$ Thus there is a directed path from each vertex $[v]\in W_{1}(\Gamma)$ to the trivial vertex corresponding to multiplying $v$ by the inverses of the $\alpha_{i}$ from $i=m$ to $i=1$.
Now suppose $W_{1}(\Gamma)$ is connected. Then there is an undirected path from every vertex $v\in\Gamma$ to the trivial loop. Observe that for every edge $[\beta]\in W_{1}(\Gamma)$ connecting $[v]$ to $[w]$, there is an edge $[\overline{\beta}]$ from $[w]$ to $[v]$. Therefore, there is a directed path in $W_{1}(\Gamma)$ from the trivial vertex to every vertex in the cover, and a directed path from the trivial loop to a vertex $[w]$ is an expression for $w$ as a product of commutators of elements of $\Gamma^{\mathrm{sep}}\cup[\pi_{1}(\Gamma),\pi_{1}(\Gamma)]$. Connectedness of $W_{1}(\Gamma)$ implies there exists such a path for every vertex in $W_{1}(\Gamma)$, so $\mathrm{H}_{1}^{\text{sep}}(\Gamma;\mathbb{Z})=\mathrm{H}_{1}(\Gamma;\mathbb{Z}).$
While we do not know whether $\mathrm{W}_{1}(\Gamma)$ is always connected, we do know that $\mathrm{W}_{1}(\Gamma)$ always has finitely many components. This follows from work of Boggi, Putman, and Salter in [BPS23].
Theorem 5.2 (Boggi–Putman–Salter).
Let $\pi:\widetilde{\Sigma}\rightarrow\Sigma$ be a finite branched covering between closed oriented surfaces. Then $\mathrm{H}_{1}^{\mathrm{pant}}(\widetilde{\Sigma};\mathbb{Q})=\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Q})$.
Here $\mathrm{H}_{1}^{\mathrm{pant}}(\widetilde{\Sigma};\mathbb{Q})$ is the subspace of $\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Q})$ spanned by
$\mathcal{C}^{\mathrm{pant}}:=\{[w]\in\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Q})\mid\pi(w)\text{ is contained in a pair of pants in }\Sigma\}.$
As the authors state in Remark 1.5 of [BPS23], this result also applies to punctured $\Sigma$ of finite type, as if $\widetilde{\Sigma}\rightarrow\Sigma$ is a finite branched cover of punctured surfaces, forgetting the punctures of $\Sigma$ and $\widetilde{\Sigma}$ yields a branched cover closed surfaces. Forgetting the punctures kills exactly the homology classes surrounding punctures, and curves surrounding punctures are simple. Since $\mathcal{C}^{\mathrm{pant}}$ contains all simple closed curves, it follows that $\mathrm{H}_{1}^{\mathrm{pant}}(\widetilde{\Sigma};\mathbb{Q})=\mathrm{H}_{1}(\widetilde{\Sigma};\mathbb{Q})$. We now show that $\mathrm{W}_{1}(\Gamma)$ has finitely many components.
Theorem 5.3.
Let $\Gamma\rightarrow\mathbf{R}_{n}$ be a finite regular cover. Then $\mathrm{W}_{1}(\Gamma)$ has finitely many components.
Proof.
Let $\mathrm{H}_{1}(\Gamma;\mathbb{Z})$ be generated by the set
$\{[\alpha_{1}],[\alpha_{2}],\ldots,[\alpha_{l}],[\beta_{1}],[\beta_{2}],\ldots,[\beta_{m}]\}$
where the $[\alpha_{i}]$ are separable homology classes, the $[\beta_{i}]$ are inseparable homology classes, and $m$ is minimal amongst generating sets for $\mathrm{H}_{1}(\Gamma;\mathbb{Z}).$ Since $\mathcal{C}^{\mathrm{pant}}\subset\mathcal{C}^{\mathrm{sep}}$, $\mathrm{H}_{1}^{\mathrm{sep}}(\Gamma;\mathbb{Z})$ is finite-index inside $\mathrm{H}_{1}(\Gamma;\mathbb{Z})$, and since $\mathrm{H}_{1}(\Gamma;\mathbb{Z})$ is abelian, every element $[\gamma]\in\mathrm{H}_{1}(\Gamma;\mathbb{Z})$ can be written as a product
$[\gamma]=[\alpha_{1}^{j_{1}}][\alpha_{2}^{j_{2}}]\ldots[\alpha_{l}^{j_{l}}][\beta_{1}^{k_{1}}][\beta_{2}^{k_{2}}]\ldots[\beta_{m}^{k_{m}}]$
where the $j_{i}$ are integers and the $k_{i}$ are nonnegative integers bounded above by $K:=[\mathrm{H}_{1}(\Gamma;\mathbb{Z}):\mathrm{H}_{1}^{\mathrm{sep}}(\Gamma;\mathbb{Z})]<\infty$. Thus there are at most $K2^{m}$ components of $\mathrm{Wh}(\Gamma)$, each corresponding to a subset of the set
$\{[\beta_{1}],[\beta_{1}^{2}],\ldots[\beta_{1}^{K-1}],[\beta_{2}],[\beta_{2}^{2}],\ldots,[\beta_{2}^{K-1}],\ldots,[\beta_{m}],[\beta_{m}^{2}],\ldots,[\beta_{m}^{K-1}]\}$
containing at most one element of $\{[\beta_{i}],[\beta_{i}^{2}],\ldots,[\beta_{i}^{K-1}]\}$ for each $i$. ∎